### v1

# Filtration learning in exact multi-parameter persistent homology and classification of time-series data

Abstract.
To analyze the topological properties of the given discrete data, one needs to consider a continuous transform called filtration. Persistent homology serves as a tool to track changes of homology in the filtration. The outcome of the topological analysis of data varies depending on the choice of filtration, making the selection of filtration crucial. Filtration learning is an attempt to find an optimal filtration that minimizes the loss function. Exact Multi-parameter Persistent Homology (EMPH) has been recently proposed, particularly for topological time-series analysis, that utilizes the exact formula of rank invariant instead of calculating it. In this paper, we propose a framework for filtration learning of EMPH. We formulate an optimization problem and propose an algorithm for solving the problem. We then apply the proposed algorithm to several classification problems. Particularly, we derive the exact formula of the gradient of the loss function with respect to the filtration parameter, which makes it possible to directly update the filtration without using automatic differentiation, significantly enhancing the learning process.
Key words and phrases:
Filtration learning, Topological time-series analysis, Optimization problem
1991 Mathematics Subject Classification:
37M10, 55N31, 65K10
1. Introduction
Topological Data Analysis (TDA) is an analysis method that infers the inherent shape of data from sampled data. Sampled data typically given discrete, so it is necessary to transform the given discrete data into continuous data, as shown in Figure 1. This transform is referred to as filtration. Persistent homology is a primary tool of TDA, utilizing barcode and persistence diagram that capture the homology changes during the filtration.
Figure 1. Offset filtration of the point cloud. The figure shows how the given discrete data are connected to form a continuous object through the filtration.
Topological data analysis and machine learning
Over the past decade, there has been continuous progress in integrating TDA and Machine Learning (ML). This attempt is composed of three parts: 1) The first step is, given a data set, to consider a filtration (e.g., Vietoris-Rips complex, cubical complex, DTM filtration, etc.) and extract a barcode, which is a summary of topological and geometric features. 2) The second step is transforming the barcode into vectors (e.g., persistence image, persistence landscape, etc.), which are suitable input data for ML. 3) The last step is inputting the transformed vectors into the ML model for training.
Figure 2. The figure shows the three steps of integrating of TDA and ML from the given raw data (left). The blue arrow in the figure indicates each step of the integration of TDA and ML.
Step 1: Since topological and geometrical analysis of discrete data itself is not plausible because they are intrinsically designed for continuous objects, it is necessary to transform the given discrete data into continuous data, as shown in Figure 1. This transformation is referred to as filtration. Therefore, extracting the topological information of the inherent shape of data depends on the choice of filtration. Commonly used filtrations include the Vietoris-Rips complex, alpha complex, and cubical complex.
Step 2: The main outcome of persistent homology is represented by barcode, but the space of barcodes lacks a natural vector space structure. This hinders the integration of TDA with many ML techniques. For example, support vector machine is based on the inner product structure, but the barcode itself is not compatible with such a structure. Therefore, we need to rearrange or transform the barcode into a vector form in a Hilbert space. This process is called the vectorization of barcodes. This has been an active research subject of numerous studies over the last decade. The persistence landscape [1] and persistence image [2] are among the well-known methods for the vectorization of barcodes.
Step 3: This step deals with the main procedure in ML. In this paper, we propose a consistent algorithm of TDA combined with ML that minimizes a task-dependent loss function using the transformed vector calculated in Step 2.
Filtration learning
Filtration learning is a process in which ML algorithms are involved in Step 1. For this, it’s necessary for Steps 1, 2, and 3 as a function to be smooth, at least locally Lipschitz. And it is also important to ensure convergence when updating each process using backpropagation. One challenge is that the space of barcodes lacks a natural smooth structure. In [3, 4, 5], it is shown that filtration learning is a smooth process under restricted conditions and it can be applied to ML problems. [6] proposes a smooth structure on the barcode space and the former studies can be explained within this framework. In [7], the authors demonstrated the convergence of stochastic gradient descent for loss functions based on persistent homology. In [8], the authors considered on filtration learning for the weight of the point in a point cloud and proposed an isometry-invariant model.
Filtration learning is similar to Convolutional Neural Network (CNN). Prior to the advent of CNNs, the filters used for image blurring or sharpening were primarily based on human experience and perception. However, with the CNN, the filter design has been automated, allowing the system to learn and adapt its own filters based on the data [9, 10]. Similarly, before the filtration learning was introduced in the realm of TDA, the design of filtrations was based on human understanding and perception. Filtration learning is to automatically design the appropriate filtration for data analysis. The methodology of TDA, including Steps 1, 2, and 3, requires heavy computation, as filtration learning necessitates that all these steps be recalculated at every epoch of backpropagation.
Exact Multi-parameter Persistent Homology
Topological time-series analysis differs from traditional statistical time-series analysis. The assumption of topological time-series analysis is that the time-series data are the observations from a dynamical system, and that involves analyzing the time-series data by examining the topological information of the dynamical system [11, 12]. This assumption differs from the one in statistics. In statistics, time series data $\left\{z_{t}\right\}_{t\in\mathbb{Z}_{\geq 0}}$ is a realized value of a certain random variable $\left\{X_{t}\right\}_{t\in\mathbb{Z}_{\geq 0}}$ for each time step [13]. Exact Multi-parameter Persistent Homology (EMPH) is one of the methods in topological time-series analysis [14]. Given time-series data, EMPH considers Fourier decomposition and regard each Fourier mode as an uncoupled harmonic oscillator, proposing the analysis based on the Liouville torus. The advantage of the proposed method is its ability to provide the exact formula of persistent homology. Traditional topological time-series analyses, such as the one using the sliding window embedding, are computationally expensive and the resulting topological measures are difficult to interpret. However, EMPH equipped with the derived exact formula, calculates persistent homology promptly. Furthermore, by constructing multi-parameter persistent homology for the Liouville torus, we can explore various topological inferences through the selection of one-parameter filtration within a multi-parameter space. In EMPH, the choice of a ray or curve in the multi-parameter space can lead to various topological inferences.
Contributions
In this paper, we propose an optimization problem for filtration learning in EMPH. We provide a detailed algorithm for filtration learning and a proof of its convergence.
Our main contribution can be summarized as follows:
  1. (1)
    We demonstrate the existence of a filtration curve that minimizes the loss function in the multi-parameter space (Theorem 3.4 in Section 3.1.1). Also, we formulate a constrained optimization problem to find such a filtration curve and provide the algorithm ((18) in Section 3.2 and Algorithm 1 in Section 4).
  2. (2)
    We derive the exact gradient formula of the loss function with respect to filtration parameters (Corollary 3.15.2 in Section 3.2). This allows us to update the filtration directly without using automatic differentiation, thereby accelerating backpropagation in filtration learning. Its efficiency is demonstrated in Example 4.1 in Section 4.
  3. (3)
    The vectorization of multi-parameter persistent homology focuses on multiple rays in a multi-parameter space, whereas our method focuses on an optimal curve in when constructing EMPH (see Remark 3.16 in Section 3.2). This approach can be generalized to an $n$-parameter space ($n\geq 2$) in a calculable manner and may yield interpretable results. Additionally, we have applied filtration learning to synthetic data and achieved improved accuracy (Examples 4.2 and 4.3). We also compared classification accuracy on a benchmark dataset in Example 4.4. With this experiment, the effectiveness of finding a filtration curve in the multi-parameter space is demonstrated. Additionally, we provide the example of the multi-parameter persistence image and landscape in Appendix C.
Outline
Section 2 reviews the basics of topological time-series analysis, EMPH theory, a smooth structure in the space of barcodes and the persistence image. Section 3, proposes an optimization problem aimed at finding an optimized curve within the EMPH framework. We compare the multi-parameter persistence image and landscape with our method. In Section 4, to explain our method effectively, we applied our theory to synthetic and benchmarking data. Section 5 provides a concluding remark and future work.
2. Background
2.1. Basic topological time-series analysis
Basic assumption in topological time-series analysis is that time-series data $\left\{z_{n}\right\}_{n\in\mathbb{Z}_{\geq 0}}$ arises from a dynamical system $(\mathcal{M},\phi)$. Here, $\mathcal{M}$ is a smooth manifold that encompasses all possible states, $\phi:\mathcal{M}\rightarrow\mathcal{M}$ is a diffeomorphism describing the state dynamics, and $y:\mathcal{M}\rightarrow\mathbb{R}$ is a smooth function representing measurements. In this setting, our assumption is summarized as follows: $z_{n}$ is expressed by $y(\phi^{n}(z_{0}))$ [11, 12]. Under this assumption, various characteristics of the dynamical system are reflected in the time-series data. Topological time-series analysis examines the time-series data by focusing on topological features of the dynamical system.
The traditional approach to time-series data analysis in TDA is to translate time-series data into a point cloud via sliding window embedding and then apply the Vietoris-Rips complex and persistent homology (see Figure 3). A periodicity score, which quantifies the periodicity of the time-series data, is proposed in [15]. This approach utilizes one-dimensional persistence information (the duration of homology in persistent homology) from the trajectory in a dynamical system.
Figure 3. Standard approach of time-series analysis in TDA.
Definition 2.1 (Sliding window embedding, [15, 16]).
For a (periodic) time-series data $f:\mathbb{R}/2\pi\mathbb{Z}:\rightarrow\mathbb{R}$, sliding window embedding $SW_{D,\tau}f$ of $f$ is a point cloud in $\mathbb{R}^{D+1}$ defined by
$SW_{D,\tau}f:=\left\{\begin{bmatrix}f(t)\\ f(t+\tau)\\ \vdots\\ f(t+D\tau)\end{bmatrix}\in\mathbb{R}^{D+1}:t\in\mathbb{R}/2\pi\mathbb{Z}\right\}.$
Definition 2.2 (Vietoris-Rips complex, [17]).
Given a metric space $(P,d)$, the Vietoris-Rips complex of $(P,d)$ is a filtration of simplicial complexes $\mathcal{R}(P)=\left\{\mathcal{R}_{\epsilon}(P)\right\}_{\epsilon\in\mathbb{R}}$, where $\mathcal{R}_{\epsilon}(P):=\bigl{\{}\left\{x_{0},\cdots,x_{n}\right\}\subset P:\bigr{.}$$\bigl{.}\max\limits_{0\leq i,j\leq n}d(x_{i},x_{j})<\epsilon\bigr{\}}$.
Sliding window embedding is motivated by Takens’ embedding theorem [11]. This mapping translates time-series data into a point cloud and enables us to apply the Vietoris-Rips complex and obtain a barcode. Given the point cloud, we infer the shape of the sampled point cloud. The Vietoris-Rips complex is a popular one-parameter filtration. A one-parameter filtration $\mathcal{X}=\left\{X_{i}\right\}_{i\in\mathbb{R}}$ is a family of topological spaces with $X_{i}\subseteq X_{j}$ for $i\leq j\in\mathbb{R}$.
Proposition 2.3 ([18]).
Let $P$ be a sampled point cloud on a closed Riemannian manifold $\mathcal{M}$, with the Gromov-Hausdorff distance between $P$ and $\mathcal{M}$ being sufficiently close. Then, for a suitable $\epsilon>0$, $\mathcal{R}_{\epsilon}(P)$ is homotopic to $\mathcal{M}$.
Proposition 2.3 provides the justification of using the point cloud $P$ to understand $\mathcal{M}$, but we do not know a priori how to pick such a suitable scale $\epsilon$. Therefore, we consider varying $\epsilon$ and calculate the homology of $\mathcal{R}_{\epsilon}(P)$. This procedure is known as persistent homology.
Definition 2.4 (Persistent homology, [19]).
Given a filtration $\mathcal{X}=\left\{X_{i}\right\}_{i\in\mathbb{R}}$, the collection of the $p$-th dimensional homology $\left\{H_{p}(X_{i})\right\}_{i\in\mathbb{R}}$ is called the $p$-th dimensional persistent homology of $\mathcal{X}$.
Multi-parameter persistence theory was introduced to address the shortcomings of single-parameter persistence theory. For example, persistent homology of the Vietoris-Rips complex is sensitive to outliers. This problem can be rectified by adding one more filtration that reflects density [20, 21]. To use multiple filtrations simultaneously, multi-parameter persistence theory is required [22]. Persistence module is a generalization of persistent homology. Now we define $n$-parameter persistence module.
Definition 2.5 ($n$-parameter persistence module [21]).
Let $(\mathbb{R}^{n},\leq)$ be an order category, where $(a_{1},\cdots,a_{n})\leq(b_{1},\cdots,b_{n})\iff a_{i}\leq b_{i}$ for all $i$, and let $\mathbf{Vect}_{\mathbb{F}}$ be the category of finite dimensional vector spaces over the field $\mathbb{F}$. The $n$-parameter persistence module is defined by a functor $\mathsf{MP}:\mathbb{R}^{n}\rightarrow\mathbf{Vect}_{\mathbb{F}}$. Persistent homology is a special case of a persistence module, which requires a filtration for its construction.
The challenge in multi-parameter persistence theory lies in the absence of a complete discrete invariant, which would determine isomorphism classes and represent the characteristics of an object in discrete values, summarizing multi-parameter persistence modules [20]. As an alternative, we consider an incomplete but practical invariant known as the fibered barcode. This approach involves one-dimensional reductions of multi-parameter persistence modules and is equivalent to the rank invariant [21].
Definition 2.6 (Barcode, [20]).
Let $\mathsf{MP}$ be a one-parameter persistence module. Then $\mathsf{MP}\cong\bigoplus\limits_{j}I(b_{j},d_{j}]$, where $I(b_{j},d_{j}]:\mathbb{R}\rightarrow\mathbf{Vect}_{\mathbb{F}}$ defined by $a\mapsto\mathbb{F}$ if $a\in(b_{j},d_{j}]$ and $a\mapsto\left\{0\right\}$ otherwise. Barcode of $\mathsf{MP}$ is defined by the collection of such $(b_{j},d_{j}]$. It is a complete invariant in the one-parameter persistence module category.
A barcode is a summary of persistent homology that simplifies our analysis. It is a complete invariant in the one-parameter persistent module category, meaning that it determines the isomorphism class.
Definition 2.7 (Fibered barcode, [21]).
Let $\mathcal{L}$ be a collection of affine lines within $\mathbb{R}^{n}$, each with a nonnegative slope. For any given line $\ell$ in $\mathcal{L}$, we consider a restriction $\mathsf{MP}^{\ell}:\ell\to\mathbf{Vect}_{\mathbb{F}}$. Since this is a one-parameter filtration, we can consider the barcode of this filtration. We refer to $\left\{\mathsf{bcd}_{*}(\mathsf{MP}^{\ell})\right\}_{\ell\in\mathcal{L}}$ as the fibered barcode of $\mathsf{MP}$.
Definition 2.8 (Rank invariant, [20, 21]).
Given an $n$-parameter persistence module $\mathsf{MP}$, the rank invariant of $\mathsf{MP}$ is defined as a function $rank(\mathsf{MP}):\left\{(s,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}:s\leq t\right\}\rightarrow\mathbb{N}$, which maps $(s,t)\mapsto rank\left(\mathsf{MP}(s)\rightarrow\mathsf{MP}(t)\right)$.
2.2. Exact Multi-parameter Persistent Homology
Recall that the assumption of topological time-series analysis is that time-series data arises from a dynamical system. Therefore, it is natural to explore which properties of time-series data can be further analyzed for specific dynamical systems with more complex structures. A Hamiltonian dynamical system $(\mathcal{M},\phi_{H})$ is a special case of dynamical system in which the trajectory of the state is governed by the Hamiltonian equation. Here, $\mathcal{M}$ is a symplectic manifold and $\phi_{H}:\mathcal{M}\rightarrow\mathcal{M}$ represents the trajectory, called a Hamiltonian diffeomorphism. EMPH considers the Liouville torus of time-series data under the Hamiltonian system of uncoupled one-dimensional harmonic oscillators. It involves constructing multi-parameter persistent homology for the Liouville torus via a multi-parameter sublevel filtration for each Fourier mode and extracting information through rays in the multi-parameter space. The EMPH consists of the following pipeline [14]:
$\begin{array}[]{ccccc}f&\longrightarrow&\mathcal{R}_{\epsilon_{1}}(r_{1}^{f}\cdot\mathbb{S}^{1})\times\cdots\times\mathcal{R}_{\epsilon_{N}}(r_{N}^{f}\cdot\mathbb{S}^{1})&\longrightarrow&\mathsf{bcd}_{n}^{\mathcal{R},\ell}(\Psi_{f})\\ \text{(Time-series data)}&&\begin{array}[]{c}\text{Multi-parameter filtration}\\ \text{of the Liouville torus}\end{array}&&\text{(Barcode)}\end{array}$
Unlike the traditional sliding window embedding method, this approach allows us to derive an exact barcode formula, providing calculable and interpretable inference of the given time-series data.
Definition 2.9 (Liouville torus of time-series data, [14]).
Given time-series data $f:\mathbb{R}/2\pi\mathbb{Z}\rightarrow\mathbb{R}$, we define the ($N$-truncated) Liouville torus $\Psi_{f,N}$ of $f$ as $\Psi_{f,N}:=r_{1}^{f}\cdot\mathbb{S}^{1}\times\cdots\times r_{N}^{f}\cdot\mathbb{S}^{1}$, where $r_{L}^{f}:=2\left|\hat{f}(L)\right|$. Here, $\hat{f}$ denotes the Fourier coefficient of $f$ and $\mathbb{S}^{1}$ is a unit circle. This torus can be represented by the Liouville torus in the system of uncoupled one-dimensional $N$ harmonic oscillators. For the details refer to [14]. From now on, unless otherwise specified, we will abbreviate $\Psi_{f,N}$ as $\Psi_{f}$.
Now, we construct multi-parameter persistent homology through a multi-parameter sublevel filtration for each Fourier mode. For each point $(\epsilon_{1},\cdots,\epsilon_{N})\in\mathbb{R}^{N}$ in the multi-parameter space, we designate the product of Vietoris-Rips complexes $\mathcal{R}_{\epsilon_{1}}\left(r_{1}^{f}\cdot\mathbb{S}^{1}\right)\times\cdots\times\mathcal{R}_{\epsilon_{N}}\left(r_{N}^{f}\cdot\mathbb{S}^{1}\right)$. Then, we consider the one-parameter reduction of the constructed multi-parameter persistent homology along a ray $\ell$. Since this reduction involves a one-parameter filtration, we can define the barcode within this one-parameter filtration. i.e., we consider the rank invariant of the multi-parameter persistent homology.
Definition 2.10.
[14] Define a multi-parameter filtration $\mathcal{R}\left(\Psi_{f}\right):\mathbb{R}^{N}\rightarrow\mathbf{Simp}$ by $\boldsymbol{\epsilon}=(\epsilon_{1},\cdots,\epsilon_{N})\mapsto\mathcal{R}_{\epsilon_{1}}\left(r_{1}^{f}\cdot\mathbb{S}^{1}\right)\times\cdots\times\mathcal{R}_{\epsilon_{N}}\left(r_{N}^{f}\cdot\mathbb{S}^{1}\right),$ where $\mathbf{Simp}$ is the category of simplicial complexes. Now, let a ray be $\ell(t)=\sqrt{N}t\cdot{\mathbf{a}\over\lVert\mathbf{a}\rVert}$ in the multi-parameter space, and define a one-parameter reduced filtration $\mathcal{R}^{\ell}\left(\Psi_{f}\right):\mathbb{R}\rightarrow\mathbf{Simp}$ specified by $\mathcal{R}^{\ell}_{t}\left(\Psi_{f}\right)=\mathcal{R}_{\sqrt{N}t\cdot{a_{1}\over\lVert\mathbf{a}\rVert}}\left(r_{1}^{f}\cdot\mathbb{S}^{1}\right)\times\cdots\times\mathcal{R}_{\sqrt{N}t\cdot{a_{N}\over\lVert\mathbf{a}\rVert}}\left(r_{N}^{f}\cdot\mathbb{S}^{1}\right)$. We refer to the barcode $\mathsf{bcd}_{*}^{\mathcal{R},\ell}(\Psi_{f})$ of this filtration as the EMPH of $f$.
Example 2.11.
For a time-series data $f(t)=\cos t+{1\over 2}\cos 2t$, its Liouville torus is $\Psi_{f}=\mathbb{S}^{1}\times{1\over 2}\cdot\mathbb{S}^{1}$. We assign each simplicial complex $\mathcal{R}_{\epsilon_{1}}(\mathbb{S}^{1})\times\mathcal{R}_{\epsilon_{2}}({1\over 2}\cdot\mathbb{S}^{1})$ to a point $(\epsilon_{1},\epsilon_{2})\in\mathbb{R}^{2}$ in the multi-parameter filtration. Figure 4 serves as a supplement to this example.
Figure 4. Multi-parameter filtration defined in Example 2.11.
Theorem 2.12.
[14] If a ray $\ell$ in the filtration space has the direction vector $\mathbf{a}=(a_{1},\cdots,a_{N})>0$, then the barcode is given by the following:
(1) $\mathsf{bcd}_{n}^{\mathcal{R},\ell}(\Psi_{f})=\left\{J_{1}^{n_{1},\ell}\bigcap\cdots\bigcap J_{N}^{n_{N},\ell}:J_{L}^{n_{L},\ell}\in\mathsf{bcd}_{n_{L}}^{\mathcal{R},\ell}\left(r_{L}^{f}\cdot\mathbb{S}^{1}\right)\ \text{and}\ \sum\limits_{L=1}^{N}n_{L}=n\right\},$
i.e.
$J_{L}^{n_{L},\ell}=\begin{cases}\left(0,\infty\right),&\mbox{if }n_{L}=0\\ \left({2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\over\sqrt{N}a_{L}/\lVert\mathbf{a}\rVert},{2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\over\sqrt{N}a_{L}/\lVert\mathbf{a}\rVert}\right],&\mbox{if }n_{L}=2k+1\\ \ \emptyset,&\mbox{otherwise}\end{cases}.$
With this theorem, we can discern which information about time-series data is encoded in the barcode of EMPH. Table 1 shows the information of time-series data encoded in the barcode.
Table 1. Relationship between EMPH and time-series data.
In [14], the authors utilized EMPH within the contexts of both unsupervised and supervised learning problems. For the unsupervised learning, it was demonstrated that the choice of ray provides a variety of perspectives for inferences. In this paper, our goal is to develop a method of finding an optimal ray. We utilize the gradient descent method to minimize the loss function and update the ray. To achieve this, it is essential to introduce a smoothness of the update process.
2.3. Smooth structure in the space of Barcodes
The motivation for updating the filtration in persistent homology, rather than adhering to a fixed filtration, is to more accurately capture the characteristics of the data within the filtration process. Therefore, we assign parameters to a collection of one-parameter filtrations and attempt to find a filtration which minimizes the loss function. The challenging aspect is that persistent homology is summarized by a barcode, but the space of barcodes lacks a natural smooth manifold structure. This difficulty hinders the application of gradient descent. Diffeology is a conceptual generalization of smooth charts that enables us to define smoothness in the space of barcodes. Motivated by diffeology theory, the following definitions are introduced in [6].
Definition 2.13.
[6] Consider smooth manifolds $\mathcal{M}$ and $\mathcal{N}$, and denote the space of barcodes by $\mathsf{Bar}$. For a map $B:\mathcal{M}\rightarrow\mathsf{Bar}$, we define it as $C^{k}$-differentiable at $x\in\mathcal{M}$ if there exists a neighborhood $U$ of $x$, $p,q\in\mathbb{Z}_{\geq 0}$ and a $C^{k}$-differentiable map $\tilde{B}:\mathcal{M}\rightarrow\mathbb{R}^{2p}\times\mathbb{R}^{q}$ such that $B=Q_{p,q}\circ\tilde{B}$ on $U$. Here, $p$ and $q$ represent the number of finite and infinite bars, respectively, and $Q_{p,q}(b_{1},d_{1},\cdots,b_{p},d_{p},v_{1},\cdots,v_{q})=\left\{(b_{1},d_{1}],\cdots,(b_{p},d_{p}],\cdots,(v_{1},\infty),\cdots,(v_{q},\infty)]\right\}$. It is important to note that both spaces, $\mathcal{M}$ and $\mathbb{R}^{2p}\times\mathbb{R}^{q}$, are smooth manifolds, each equipped with a natural smooth structure. A map $V:\mathsf{Bar}\rightarrow\mathcal{N}$ is called $C^{k}$-differentiable at $\mathcal{B}\in\mathsf{Bar}$ if for all $p,q\in\mathbb{Z}_{\geq 0}$ and ~$\tilde{\mathcal{B}}\in\mathbb{R}^{2p}\times\mathbb{R}^{q}$ such that $Q_{p,q}(\tilde{\mathcal{B}})=\mathcal{B}$, the map $\tilde{V}=V\circ Q_{p,q}:\mathbb{R}^{2p}\times\mathbb{R}^{q}\rightarrow\mathcal{N}$ is a $C^{k}$-differentiable map on a neighborhood of $\tilde{\mathcal{B}}$. Figure 5 is a schematic illustration of this definition.
Figure 5. Schematic illustration of smooth structure on the space of barcodes.
In machine learning, it is valid to introduce such a differential structure to the barcode space when searching for parameters to minimize loss functions. What we desire is that when $V$ and $B$ are smooth defined in Definition 2.13, $V\circ B:\mathcal{M}\rightarrow\mathcal{N}$ is smooth within the smooth manifold category.
Proposition 2.14.
[6] Suppose $B:\mathcal{M}\rightarrow\mathsf{Bar}$ and $V:\mathsf{Bar}\rightarrow\mathcal{N}$ are $C^{k}$-differentiable. Then $V\circ B:\mathcal{M}\rightarrow\mathcal{N}$ is a well-defined (the derivative of $V\circ B$ is independent of $\tilde{B}$ and $\tilde{D}$) $C^{k}$-differentiable function in the smooth manifold category.
2.4. Persistence Image
Note that the space of barcodes lacks not only a natural differential structure but also a vector space structure. This limitation hinders the integration of TDA with ML techniques. However, several studies have been conducted to develop mappings from barcodes to a Hilbert space; among theses, the persistence landscape [1] and persistence image [2] are well-known methods. In this study, we use the vectorization with the persistence image because it has been demonstrated to be a smooth map in [6], suitable for analyzing the smoothness of our workflow.
Definition 2.15.
[2] Given a barcode $\mathcal{B}$, its persistence image is an $r^{2}$-dimensional vector $\mathbf{I}_{\mathcal{B}}=\bigl{(}I_{\mathcal{B}}(x_{1},y_{1}),\cdots,$ $I_{\mathcal{B}}(x_{r^{2}},y_{r^{2}})\bigr{)}\in\mathbb{R}^{r^{2}}$, where
(2) $I_{\mathcal{B}}(x,y)=\sum\limits_{(b_{i},d_{i}]\in\mathcal{B}}\omega(d_{i}-b_{i})g_{b_{i},d_{i}}(x,y),$
and $\omega:\mathbb{R}\rightarrow\mathbb{R}$ is a weight function and $g_{b_{i},d_{i}}(x,y)$ is defined as ${1\over 2\pi\sigma^{2}}e^{-{{\left[(x-b_{i})^{2}+(y-(d_{i}-b_{i}))^{2}\right]}\over 2\sigma^{2}}}$. Here, $(x_{j},y_{j})$ represents a subdivided point within the given square. From now on, unless otherwise specified, we will abbreviate $\mathbf{I}_{\mathcal{B}}$ as $\mathbf{I}$.
In [2], the persistence image is defined as the integration of $g_{b_{i},d_{i}}$. However, for practical purpose, the above discrete form is usualy used, e.g. in Gudhi [23]. The following propositions highlight the advantageous properties of the persistence image.
Proposition 2.16 ([2]).
Let $W_{1}$ be the $1$-Wasserstein distance on the space of barcodes. The persistence image $\mathbf{I}:(\mathsf{Bar},W_{1})\rightarrow\mathbb{R}^{n}$ is a Lipschitz function.
Given that for barcodes $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$, $W_{p}(\mathcal{B}_{1},\mathcal{B}_{2})\leq W_{q}(\mathcal{B}_{1},\mathcal{B}_{2})$ when $p\geq q$, it follows that Proposition 2.16 also holds for the bottleneck distance $(\mathsf{Bar},d_{B})$ [24].
Proposition 2.17.
[6] If the weight function $\omega$ is $C^{k}$-differentiable, then the persistence image is $C^{k}$-differentiable.
3. Filtration learning within Exact Multi-parameter Persistent Homology
In this section, we propose a time-series classifier model that updates the filtration. Our problem can be summarized as an optimization problem, and Figure 6 schematically illustrates the proposed model. We demonstrate the existence of an optimal filtration ray, or more generally, an optimal filtration curve, in the multi-parameter space. Here, ‘optimal filtration’ means the filtration that minimizes the loss function. First, we address the problem of finding an optimal filtration ray in Section 3.1, and then we generalize problem that addresses an optimal filtration curve in Section 3.2.
3.1. Filtration ray learning
In this section, we discuss how to find an optimal filtration ray for classification problems. We refer this process to as filtration ray learning, as defined in (3). The term ‘optimal filtration ray’ refers to a filtration ray that minimizes the loss function.
Figure 6 summarizes our model. Given time-series data, we apply EMPH to obtain a barcode, then transform it into a persistence image (PI) and input this into the neural network. In the multi-parameter space, if the direction vector of a ray changes, EMPH changes, and as a result, PI changes. Consequently, the input to the neural network changes as well. In this scenario, we aim to perform backpropagation to identify a direction vector that minimizes the loss function.
Figure 6. Summary of our model. We transform time-series data into a persistence image (PI) using EMPH, input it into a neural network, and use backpropagation to find the optimal direction vector.
Now, define functions as follows in Table 2.
Table 2. Definition of functions in filtration ray learning.
Our problem can be summarized as the following constrained optimization problem:
(3) $\operatorname*{arg\,min}_{\begin{subarray}{c}\mathbf{W},\mathbf{b},\\ \mathbf{a}\in\mathbb{R}_{m,M}^{N}\end{subarray}}\mathcal{L}(\mathbf{W},\mathbf{b},\mathbf{a}),$
where $\mathbf{W}$ and $\mathbf{b}$ represent the weight and bias in the neural network, respectively, and $\mathbf{a}$ is a direction vector of a ray in the multi-parameter space which is constrained to $\mathbb{R}_{m,M}^{N}$. What is remarkable about the proposed optimization problem is that the loss function includes $\mathbf{a}$ as a variable. We want to find a minimizer $\mathbf{a}$ which minimizes the loss function. We demonstrate the smoothness of $\mathcal{L}$ and explain how to solve the optimization problem in the following section.
Remark 3.1.
We applied Min-Max scaling $\Phi$ to the persistence image, which is the input to the neural network. At each epoch, the persistence image is updated, with its scale being changed. Min-Max scaling yields better results according to our numerical experiments.
3.1.1. Smoothness of framework
In the filtration learning of EMPH, gradient descent is used to minimize the loss function, and the backpropagation algorithm is employed to update the direction vector. To achieve convergence, we first demonstrate the smoothness of our model. First, we examine the smoothness of one-dimensional EMPH because it exhibits different behavior from that of higher dimensions.
Lemma 3.2.
Let $B:\mathbb{R}_{>0}^{N}\rightarrow\mathsf{Bar}$ be a map where $\mathbf{a}\mapsto\mathsf{bcd}_{1}^{\mathcal{R},\ell}(\Psi_{f})$. Then $B$ is a smooth map.
Proof.
Define $\tilde{B}:\mathbb{R}_{>0}^{N}\rightarrow\mathbb{R}^{2N}$ by $\tilde{B}(\mathbf{a})=\left(b_{1}(\mathbf{a}),d_{1}(\mathbf{a}),b_{2}(\mathbf{a}),d_{2}(\mathbf{a}),\cdots,b_{N}(\mathbf{a}),d_{N}(\mathbf{a})\right)$, where $b_{L},d_{L}:\mathbb{R}_{>0}^{N}\rightarrow\mathbb{R}$ by $b_{L}(\mathbf{a})=0$ and $d_{L}(\mathbf{a})={\sqrt{3}r_{L}^{f}\over\sqrt{N}a_{L}/\lVert\mathbf{a}\rVert}$. Then $\tilde{B}$ is clearly smooth map for every $\mathbf{a}\in\mathbb{R}_{>0}^{N}$ and $Q_{N,0}\circ\tilde{B}(\mathbf{a})=\mathsf{bcd}_{1}^{\mathcal{R},\ell}(\Psi_{f})=B(\mathbf{a})$. ∎
However, in higher dimensions, $B$ may not be smooth, and similarly, neural networks such as ReLU networks may also not be smooth. But this is not a crucial issue in optimization problems. The important thing is that these functions are locally Lipschitz since we can consider the subgradient descent method; Lemma 3.5.
Remark 3.3.
If $n>1$, then $B:\mathbb{R}_{>0}^{N}\rightarrow\mathsf{Bar},\mathbf{a}\mapsto\mathsf{bcd}_{n}^{\mathcal{R},\ell}(\Psi_{f})$ is continuous but may not be smooth. For example, consider a ray $\ell(t)=\sqrt{N}t\cdot{\mathbf{a}\over\lVert\mathbf{a}\rVert}$. Since the minimum function is not smooth, $\mathsf{bcd}_{2}^{\mathcal{R},\ell}(\Psi_{f})=\Biggl{\{}\left(0,\min\left({r_{L_{i_{1}}}^{f}\over\sqrt{N}a_{i_{1}}/\lVert\mathbf{a}\rVert},{r_{L_{i_{2}}}^{f}\over\sqrt{N}a_{i_{2}}/\lVert\mathbf{a}\rVert}\right)\sqrt{3}\right]:1\leq i_{1}<i_{2}\leq N\Biggr{\}}$ is not smooth.
Now, we introduce one of the main theorems, i.e., the existence of a ray that minimizes the loss function.
Theorem 3.4 (Variational principle of Exact Multi-parameter Persistent Homology).
For a compact subset $\mathcal{C}\subset\mathbb{R}_{>0}^{N}$, the function $\mathcal{L}|_{\mathcal{C}}:\mathcal{C}\rightarrow\mathbb{R}$ is continuous, ensuring that it has a minimum value.
This theorem signifies that there is an optimal ray, but, (3) is not convex in general. Thus the uniqueness of the optimal ray is not guaranteed. Following this theorem, the next step is to demonstrate how we can find such a ray.
3.1.2. Convergence analysis
In the optimization problem presented in (3), it is still possible to find a feasible solver even if $\mathcal{L}$ is neither smooth nor convex which includes the situations mentioned in Remark 3.3, or when the activation function is non-smooth, such as the ReLU function commonly used in neural networks. Therefore, it is necessary to introduce a generalization of the gradient to minimize the nonsmooth, nonconvex loss function. The Clarke subgradient serves as a generalization of the gradient for nonsmooth and nonconvex functions. According to Rademacher’s theorem, a locally Lipschitz function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ has a Clarke subgradient. If $f$ is smooth, then the Clarke subgradient is the gradient of $f$. If $f$ is a convex function, then it is the subgradient of $f$ [25]. The following lemma guarantees the convergence of the projected stochastic subgradient descent method using the Clarke subgradient. From this point, we utilize the properties of definable functions to prove Theorem 3.8. In an o-minimal structure, a definable function, roughly speaking, satisfies the properties of a piecewise smooth function, and a semi-algebraic function is a special case of a definable function. The basic concepts of o-minimal structures and definable functions are summarized in Appendix A.
Figure 7. Schematic illustration of filtration ray learning of EMPH.
Lemma 3.5 (Corollary 6.4, [26]).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be both a locally Lipschitz function and a definable function. Consider the projected stochastic subgradient descent method
[…]
where $\left\{\zeta(x_{k},\omega_{k})\right\}_{k}$ is a stochastic estimator for the Clarke subgradient of $f$ at $x_{k}$, $\left\{\alpha_{k}\right\}_{k\geq 1}$ is a sequence of step-sizes, $\mathcal{X}$ is a closed and definable constraint region, and $\mathsf{proj}_{\mathcal{X}}:\mathbb{R}^{d}\rightarrow\mathcal{X}$ is a projection map. If the following conditions are satisfied, then $x_{k}$ almost surely converges to Clarke stationary point:
  1. (1)
    (Proper step-size) $\alpha_{k}\geq 0$, $\sum\limits_{k=1}^{\infty}\alpha_{k}=\infty$ and $\sum\limits_{k=1}^{\infty}\alpha_{k}^{2}<\infty$.
  2. (2)
    (Boundedness) $\sup\limits_{k\geq 1}\lVert x_{k}\rVert<\infty$
  3. (3)
    (Noise constraint) There exists a function $p:\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ that is bounded on bounded sets, such that expectation $\mathbb{E}_{\omega}\left[\zeta(x,\omega)\right]\in\partial f(x),\mathbb{E}_{\omega}\left[\lVert\zeta(x_{k},\omega_{k})\rVert^{2}\right]\leq p(x)$ for all $x\in\mathcal{X}$ and $\mathbb{E}_{\omega}\left[\sup\limits_{k\geq 1}\lVert\zeta(z_{k},\omega)\rVert\right]<\infty$ for every convergent sequence $\left\{z_{k}\right\}_{k\geq 1}$.
To apply Lemma 3.5 to (3) in our case, we need to prove that $\mathcal{L}$ is both locally Lipschitz and definable. Our strategy involves verifying that both $\tilde{B}$ and $\tilde{V}$ are locally Lipschitz and definable functions, which leads to $\mathcal{L}$ being locally Lipschitz and definable. In this section, the main theorem is Theorem 3.8. To prove this, Lemma 3.6 demonstrates that $\tilde{B}$ is locally Lipschitz and definable. Lemma 3.7 shows that $\tilde{V}$ is locllay Lipschitz and definable.
Here, we need to understand the properties of definable functions to prove Theorem 3.8. The properties of definable function are detailed in Appendix A, and minor propositions regarding locally Lipschitz functions along with their proofs are in Appendix B.
Lemma 3.6.
Let $\tilde{B}:\mathbb{R}^{N}_{>0}\rightarrow\mathbb{R}^{2p}\times\mathbb{R}^{q}$ be a map that satisfies $B=Q_{p,q}\circ\tilde{B}$. Then $\tilde{B}$ is locally Lipschitz and definable.
Proof.
Let
(4) $b_{L,n_{L}}(\mathbf{a})=\begin{cases}{2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\over\sqrt{N}a_{L}/\lVert\mathbf{a}\rVert},&\mbox{if }n_{L}=2k+1\\ 0,&\mbox{otherwise }\end{cases},$
(5) $d_{L,n_{L}}(\mathbf{a})=\begin{cases}{2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\over\sqrt{N}a_{L}/\lVert\mathbf{a}\rVert},&\mbox{if }n_{L}=2k+1\\ \infty,&\mbox{otherwise }\end{cases}.$
Note that for any $(b,d]\in\mathsf{bcd}_{n}^{\mathcal{R},\ell}(\Psi_{f})$, $b$ is expressed by the maximum of $\left\{b_{L,n_{L}}\right\}_{L}$ and $d$ is expressed by the minimum of $\left\{d_{L,n_{L}}\right\}_{L}$ in Theorem 2.12. Since $b_{L,n_{L}}$ and $d_{L,n_{L}}$ are smooth functions on a compact subset of $\mathbb{R}_{>0}^{N}$, by Proposition B.2, they are locally Lipschitz. By Proposition B.3, both $b$ and $d$ are locally Lipschitz. Similarly, the maps $b_{L,n_{L}}$ and $d_{L,n_{L}}$ are semialgebraic by Proposition A.6, and their maximum and minimum are also semialgebraic by Proposition A.7. Therefore, by Proposition A.8, $\tilde{B}$ is definable. ∎
Lemma 3.7.
$\tilde{V}=l(y,\cdot)\circ\phi\circ NN\circ\Phi\circ\mathbf{I}\circ Q_{p,q}$ is locally Lipschitz and definable.
Proof.
  1. (1)
    $l(y,\cdot)\circ\phi$ is smooth function on a compact set, by Proposition B.2, it is locally Lipschitz.
  2. (2)
    $NN$ is locally Lipschitz since it consists of affine transformations and ReLU functions.
  3. (3)
    $\mathbf{I}$ is locally Lipschitz, as established by Proposition 2.16.
  4. (4)
    $Q_{p,q}$ is locally Lipschitz by Proposition 3.2 in [6].
  5. (5)
    $\mathbf{I}\circ Q_{p,q}$, ReLU, softmax, and cross-entropy functions are definable functions.
[…]
  6. (6)
    Claerly, $\Phi$ is locally Lipschitz and definable.
Since every component of $\tilde{V}$ consists of locally Lipschitz functions and definable functions, $\tilde{V}$ is locally Lipschitz and definable. ∎
For simplicity in expressing the gradient in Section 3.1.3, define $\boldsymbol{\rho}$ as $\boldsymbol{\rho}(\mathbf{a})={\mathbf{a}\over\lVert\mathbf{a}\rVert}$. Now, consider the following update rule:
(6) $W^{[l]}_{k+1}$ $=W^{[l]}_{k}-\alpha_{k}{\partial\mathcal{L}\over\partial{W^{[l]}_{k}}},$
$b^{[l]}_{k+1}$ $=b^{[l]}_{k}-\alpha_{k}{\partial\mathcal{L}\over\partial{b^{[l]}_{k}}},$
$\mathbf{a}_{k+{1\over 2}}$ $=\mathbf{a}_{k}-\alpha_{k}{\partial\mathcal{L}\over\partial{\mathbf{a}_{k}}},$
$\mathbf{a}_{k+1}$ $=\mathsf{proj}_{\mathbb{R}^{N}_{m,M}}\left(\boldsymbol{\rho}\left(\mathbf{a}_{k+{1\over 2}}\right)\right).$
Note that the problem (3) is inherently nonconvex and nonsmooth, so finding a globally optimal solution is challenging. Alternatively, our aim is to find a stationary point within the loss function. Note that Clarke stationary points need not be local minima [27].
Let $\mathbf{W}_{k}:=\left(W^{[1]}_{k},\cdots,W^{[h]}_{k}\right)$ and $\mathbf{b}_{k}:=\left(b^{[1]}_{k},\cdots,b^{[h]}_{k}\right)$. From the following theorem, we can solve the problem (3).
Theorem 3.8.
The sequences $\mathbf{W}_{k},\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ in (6) converge to Clarke stationary points of $\mathcal{L}$.
Proof.
Note that $V\circ B=V\circ Q_{p,q}\circ\tilde{B}=\tilde{V}\circ\tilde{B}$. By Lemmas 3.6, 3.7 and 3.5, this statement is true. ∎
Remark 3.9.
At each epoch, we update the direction of the rays, which changes the barcode. In filtration learning, when calculating the persistence image within a fixed grid on the birth-persistence axis (here, “persistence” refers to the difference between the birth and death times in the barcode), if the persistence of a barcode increases over epochs and moves outside the grid on the birth-persistence axis during the filtration learning, the persistence of the barcode would be decreased again, potentially inducing oscillations. This appears to interfere with the process of updating the direction (see Figure 8). Therefore, we should restrict the domain from $\mathbb{R}_{>0}^{N}$ to $\mathbb{R}_{m,M}^{N}$, where $\mathbb{R}_{m,M}=\left\{x\in\mathbb{R}:0<m\leq r\leq M\right\}$, $m=c_{1}\min\limits_{L}b_{L}(\mathbf{a})$, and $M=c_{2}\max\limits_{L}d_{L}(\mathbf{a})$. This setting maintains the stability of updating the direction vectors. In Section 4, we set $c_{1}=\frac{1}{2}$ and $c_{2}=2$.
Figure 8. If a point moves out of the grid on the birth-persistence space, its impact on the persistence image decreases. Therefore, we should limit the extent of the updated direction to ensure that all points remain within the grid.
3.1.3. Exact formula of gradient
In this section, we discuss a method for efficiently updating the filtration ray in the backward process. In Figure 6, the forward process demonstrates that the exact barcode formula allows efficient computation. The computation of the barcode involves multiple intermediate steps, which makes the procedure of updating the direction through automatic differentiation inefficient. By using the gradient of the loss function with respect to the direction vector, derived from the barcode formula, we can update the direction vector directly without relying on automatic differentiation, thereby allowing efficient computation of the backward process. In Example 4.1, that is demonstrated experimentally.
Lemma 3.10.
The derivative of $\boldsymbol{\rho}(\mathbf{a})={\mathbf{a}\over\lVert\mathbf{a}\rVert}=(\rho_{1},\cdots,\rho_{N})$ is given by the following:
${\partial\boldsymbol{\rho}\over\partial\mathbf{a}}={1\over\lVert\mathbf{a}\rVert}\mathbf{1}-{1\over\lVert\mathbf{a}\rVert^{3}}\mathbf{a}\mathbf{a}^{T}.$
Here, $\mathbf{1}$ is an identity matrix.
Proof.
The partial derivative of the $i$-th entry of $\boldsymbol{\rho}(\mathbf{a})$ with respect to the variable $a_{L}$ is as follows:
  1. (1)
    If $i=L$,
    ${\partial\rho_{i}\over\partial a_{L}}$ $={\partial\over\partial a_{L}}\left({a_{i}\over\sqrt{a_{1}^{2}+\cdots+a_{N}^{2}}}\right)$
    $=1\cdot{1\over\sqrt{a_{1}^{2}+\cdots+a_{N}^{2}}}+a_{L}\cdot-{1\over 2}\cdot{2a_{L}\over\left(\sqrt{a_{1}^{2}+\cdots+a_{N}^{2}}\right)^{3}}$
    $={1\over\lVert\mathbf{a}\rVert}-{a_{L}^{2}\over\lVert\mathbf{a}\rVert^{3}}.$
  2. (2)
    If $i\neq L$,
    ${\partial\rho_{i}\over\partial a_{L}}={\partial\over\partial a_{L}}\left({a_{i}\over\sqrt{a_{1}^{2}+\cdots+a_{N}^{2}}}\right)=-{a_{i}a_{L}\over\lVert\mathbf{a}\rVert^{3}}.$
This can be summarized in the following matrix form as:
${\partial\boldsymbol{\rho}\over\partial\mathbf{a}}={1\over\lVert\mathbf{a}\rVert}\mathbf{1}-{1\over\lVert\mathbf{a}\rVert^{3}}\mathbf{a}\mathbf{a}^{T}.$
Theorem 3.11 (Exact formula of gradients).
The explicit forms of the gradients in (6) are as follows:
(7) ${\partial\mathcal{L}\over\partial W^{[l]}}$ $=\left(\delta^{[l]}\right)\left(g^{[l-1]}(z_{l-1})\right)^{T},$
${\partial\mathcal{L}\over\partial b^{[l]}}$ $=\left(\delta^{[l]}\right),$
${\partial\mathcal{L}\over\partial\mathbf{a}}$ $=\left(\delta^{[1]}W^{[1]}\right)\left({1\over\mathbf{I}_{\max}-\mathbf{I}_{\min}}\right)\left({\partial\mathbf{I}\over\partial\mathsf{b}}{\partial\mathsf{b}\over\partial\boldsymbol{\rho}}+{\partial\mathbf{I}\over\partial\mathsf{d}}{\partial\mathsf{d}\over\partial\boldsymbol{\rho}}\right)\left({1\over\lVert\mathbf{a}\rVert}\mathbf{1}-{1\over\lVert\mathbf{a}\rVert^{3}}\mathbf{a}\mathbf{a}^{T}\right),$
where $z_{l}=W^{[l]}g^{[l-1]}(z_{l-1})+b^{[l]},\delta^{[l]}={\partial\mathcal{L}\over\partial z_{l}}$ and the persistence image $z_{0}=\mathbf{I}=(I_{1},\cdots,I_{r^{2}}),$ with $\mathbf{I}_{max}=\max(I_{1},\cdots,I_{r^{2}}),\mathbf{I}_{min}=\min(I_{1},\cdots,I_{r^{2}})$. Additionally, $\mathsf{b}=(b_{1},\cdots,b_{p})$ and $\mathsf{d}=(d_{1},\cdots,d_{p})$ represent the coordinates of birth and death times as defined in Definition 2.13, and $\boldsymbol{\rho}(\mathbf{a})={\mathbf{a}\over\lVert\mathbf{a}\rVert}=(\rho_{1},\cdots,\rho_{N})$.
Each component in ${\partial\mathcal{L}\over\partial\mathbf{a}}$ is
${\partial I_{v}\over\partial b_{j}}=$ $-\sum\limits_{(b_{i},d_{i}]\in\mathcal{B}}g_{b_{i},d_{i}}(x_{v},y_{v})$
$+$ $(d_{j}-b_{j})\cdot{d_{j}-2b_{j}\over\sigma^{2}}\sum\limits_{(b_{i},d_{i}]\in\mathcal{B}}g_{b_{i},d_{i}}(x_{v},y_{v})$
(8) $+$ $(d_{j}-b_{j})\cdot{1\over\sigma^{2}}\sum\limits_{(b_{i},d_{i}]\in\mathcal{B}}(x_{v}-y_{v})g_{b_{i},d_{i}}(x_{v},y_{v}),$
${\partial I_{v}\over\partial d_{j}}=$ $\sum\limits_{(b_{i},d_{i}]\in\mathcal{B}}g_{b_{i},d_{i}}(x_{v},y_{v})$
-- $(d_{j}-b_{j})\cdot{d_{j}-b_{j}\over\sigma^{2}}\sum\limits_{(b_{i},d_{i}]\in\mathcal{B}}g_{b_{i},d_{i}}(x_{v},y_{v})$
(9) $+$ $(d_{j}-b_{j})\cdot{1\over\sigma^{2}}\sum\limits_{(b_{i},d_{i}]\in\mathcal{B}}y_{v}g_{b_{i},d_{i}}(x_{v},y_{v}),$
and Clarke subgradients
(10) ${\partial b_{j}\over\partial\rho_{L}}\in$ $\ {\partial\over\partial\rho_{L}}\left(\max\left\{b_{L,n_{L}}(\boldsymbol{\rho}):\sum\limits_{L=1}^{N}n_{L}=n\right\}\right),$
(11) ${\partial d_{j}\over\partial\rho_{L}}\in$ $\ {\partial\over\partial\rho_{L}}\left(\min\left\{d_{L,n_{L}}(\boldsymbol{\rho}):\sum\limits_{L=1}^{N}n_{L}=n\right\}\right),$
where $(b_{j},d_{j}]=J_{1}^{n_{1},\ell}\bigcap\cdots\bigcap J_{N}^{n_{N},\ell}\in\mathsf{bcd}_{n}^{\mathcal{R},\ell}(\Psi_{f})$ in (1), and $b_{L,n_{L}}$ and $d_{L,n_{L}}$ are defined in (4) and (5), respectively. In particular, if we consider a one-dimensional barcode $(n=1)$, $j$ can be indexed such that $(b_{j},d_{j}]=\Bigl{(}0,\sqrt{{3\over N}}{r_{j}^{f}\over\rho_{j}}\Bigr{]}$. Then, we get
(12) ${\partial b_{j}\over\partial\rho_{L}}=$ $\ 0\quad\text{ for }1\leq j,L\leq N,$
(13) ${\partial d_{j}\over\partial\rho_{L}}=$$\begin{cases}-\sqrt{{3\over N}}{r_{L}^{f}\over\rho_{L}^{2}},&\mbox{if }j=L\\ 0,&\mbox{otherwise}\end{cases}.$
Proof.
The derivatives ${\partial\mathcal{L}\over\partial W^{[l]}}$ and ${\partial\mathcal{L}\over\partial b^{[l]}}$ are well-known gradients in deep neural networks [28],
${\partial\mathcal{L}\over\partial\mathbf{a}}$ $={\partial\mathcal{L}\over\partial\Phi}{\partial\Phi\over\partial\mathbf{I}}\left({\partial\mathbf{I}\over\partial\mathsf{b}}{\partial\mathsf{b}\over\partial\boldsymbol{\rho}}+{\partial\mathbf{I}\over\partial\mathsf{d}}{\partial\mathsf{d}\over\partial\boldsymbol{\rho}}\right){\partial\boldsymbol{\rho}\over\partial\mathbf{a}}$
$=\left(\delta^{[1]}W^{[1]}\right)\left({1\over\mathbf{I}_{\max}-\mathbf{I}_{\min}}\right)\left({\partial\mathbf{I}\over\partial\mathsf{b}}{\partial\mathsf{b}\over\partial\boldsymbol{\rho}}+{\partial\mathbf{I}\over\partial\mathsf{d}}{\partial\mathsf{d}\over\partial\boldsymbol{\rho}}\right)\left({1\over\lVert\mathbf{a}\rVert}\mathbf{1}-{1\over\lVert\mathbf{a}\rVert^{3}}\mathbf{a}\mathbf{a}^{T}\right)$
Equations (8) and (9) can be directly computed from the definition of $\mathbf{I}$ in (2). Equations (10) and (11) are induced from $(b_{j},d_{j}]=J_{1}^{n_{1},\ell}\bigcap\cdots\bigcap J_{N}^{n_{N},\ell}=\biggl{(}\max\Bigl{\{}b_{L,n_{L}}(\boldsymbol{\rho}):\sum\limits_{L=1}^{N}n_{L}=n\Bigr{\}},\min\Bigl{\{}d_{L,n_{L}}(\boldsymbol{\rho}):\sum\limits_{L=1}^{N}n_{L}=n\Bigr{\}}\biggr{]}$. For (12) and (13), set $b_{j}=\max\left\{b_{j,n_{j}}(\boldsymbol{\rho}^{1},\cdots,\boldsymbol{\rho}^{R}):n_{j}=1\text{ and }n_{1},\cdots,n_{j-1},n_{j+1},\cdots,n_{N}=0\right\}=0$ and $d_{j}=\min\bigl{\{}d_{j,n_{j}}(\boldsymbol{\rho}^{1},\cdots,\boldsymbol{\rho}^{R}):n_{j}=1\text{ and }n_{1},\cdots,n_{j-1},n_{j+1},\cdots,n_{N}=0\bigr{\}}=\sqrt{{3\over N}}{r_{j}^{f}\over\rho_{j}}$. Differentiating these with respect to $\rho_{L}$ yields (12) and (13). ∎
From this formula, we can directly update $\mathbf{a}$ in (6) and calculate efficiently the backward process. In Example 4.1, we compare the computation times of filtration ray learning among three cases: 1) Forward and backward processes are conducted on GPUs and direction vector $\mathbf{a}$ is updated directly using the formula of ${\partial\mathcal{L}\over\partial\mathbf{a}}$ in (7). And the neural network is updated using automatic differentiation in PyTorch. 2) Forward and backward processes are conducted on GPUs and all parameters are updated using automatic differentiation in PyTorch. 3) Forward and backward processes are conducted on CPUs and all parameters are updated using the formula (7) in Numpy + Scikit-learn.
3.2. Filtration curve learning
Until now, we have demonstrated the process of finding the optimal filtration ray within a multi-parameter space. This approach can be generalized to curved filtration, a refinement that offers a more accurate reflection of the characteristics of the data.
Definition 3.12 (Filtration curve and persistent homology).
For an increasing curve filtration $\mathbf{c}=(c_{1},\cdots,c_{N}):[0,Q]\rightarrow\mathbb{R}^{N}$ (with $\mathbf{c}^{\prime}>0$), we define a one-parameter filtration $\mathcal{R}^{\mathbf{c}}\left(\Psi_{f}\right):\mathbb{R}\rightarrow\mathbf{Simp}$ by setting $\mathcal{R}^{\mathbf{c}}_{t}\left(\Psi_{f}\right)=\mathcal{R}_{c_{1}(t)}\left(r_{1}^{f}\cdot\mathbb{S}^{1}\right)\times\cdots\times\mathcal{R}_{c_{N}(t)}\left(r_{N}^{f}\cdot\mathbb{S}^{1}\right)$. We denote the barcode induced from this filtration as $\mathsf{bcd}_{*}^{\mathcal{R},\mathbf{c}}(\Psi_{f})$.
As in the proof of Theorem 2.12, we can easily deduce the exact formula for the barcode $\mathsf{bcd}_{*}^{\mathcal{R},\mathbf{c}}(\Psi_{f})$.
Lemma 3.13.
Let an increasing curve $\mathbf{c}:[0,Q]\rightarrow\mathbb{R}^{N}$ with $\mathbf{c}(0)=\mathbf{0}$ have a constant speed of $\sqrt{N}$. Then the barcode is given by the following:
(14) $\mathsf{bcd}_{n}^{\mathcal{R},\mathbf{c}}(\Psi_{f})=\left\{J_{1}^{n_{1},\mathbf{c}}\bigcap\cdots\bigcap J_{N}^{n_{N},\mathbf{c}}:J_{L}^{n_{L},\mathbf{c}}\in\mathsf{bcd}_{n_{L}}^{\mathcal{R},\mathbf{c}}\left(r_{L}^{f}\cdot\mathbb{S}^{1}\right)\ \text{and}\ \sum\limits_{L=1}^{N}n_{L}=n\right\},$
i.e.
$J_{L}^{n_{L},\mathbf{c}}=\begin{cases}\left(0,\infty\right),&\mbox{if }n_{L}=0\\ \left(c_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\right),c_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\right)\right],&\mbox{if }n_{L}=2k+1\\ \ \emptyset,&\mbox{otherwise}\end{cases}.$
Proof.
Note that
$\mathsf{bcd}_{n}^{\mathcal{R},\mathbf{c}}(r_{L}^{f}\cdot\mathbb{S}^{1})=\begin{cases}\left\{\left(0,\infty\right)\right\},&\mbox{if }n=0\\ \left\{J_{L}^{2k+1,\mathbf{c}}\right\},&\mbox{if }n=2k+1\\ \emptyset,&\mbox{ otherwise}\end{cases},$
where
$J_{L}^{2k+1,\mathbf{c}}:=\left(\min\limits_{t}c_{L}(t)\geq 2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right),\min\limits_{t}c_{L}(t)\geq 2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\right]$
$=\left(c_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\right),c_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\right)\right].$
By applying the persistent Künneth formula ([29, 14]) to the metric space $(r_{1}^{f}\cdot\mathbb{S}^{1}\times\cdots\times r_{N}^{f}\cdot\mathbb{S}^{1})$, with maximum metric, we obtain (14).
We can easily verify that the barcode formula for filtration ray is a special case of filtration curve.
Example 3.14.
Let $c_{L}(t)=\sqrt{N}t\cdot{a_{L}\over\lVert\mathbf{a}\rVert}$. Then we have its inverse $c_{L}^{-1}(t)={t\over\sqrt{N}a_{L}/\lVert\mathbf{a}\rVert}$ and $\mathsf{bcd}_{n}^{\mathcal{R},\mathbf{c}}(\Psi_{f})=\mathsf{bcd}_{n}^{\mathcal{R},\ell}(\Psi_{f})$ as in Theorem 2.12.
Extending the filtration ray optimization directly to the filtration curves is not straightforward. This is due to the fundamental difference them. That is the classification with the filtration ray is parameterized by a single direction vector, but the filtration curve is generally parameterized continuously, not finitely unlike the filtration ray. This means that filtration curve learning involves infinitely many parameter updates. Consequently, we address this issus by approximating the curve as piecewise linear segments, which allows the filtration curve to be parameterized finitely.
Given a curve $\mathbf{c}=(c_{1},\cdots,c_{N}):[0,Q]\rightarrow\mathbb{R}^{N}$ with $\mathbf{c}(0)=\mathbf{0}$, consider a piecewise line segments approximation $\tilde{\mathbf{c}}=(\tilde{c}_{1},\cdots,\tilde{c}_{N}):[0,Q]\rightarrow\mathbb{R}^{N}$, where
(15) $\tilde{c}_{L}(t)=\sqrt{N}{Q\over R}\sum\limits_{i=1}^{s-1}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}+\sqrt{N}\left(t-{(s-1)Q\over R}\right){a_{L}^{s}\over\lVert\mathbf{a}^{s}\rVert},\quad(s-1){Q\over R}\leq t\leq s{Q\over R},$
$\mathbf{a}^{s}=(a^{s}_{1},\cdots,a^{s}_{N})>0$ and $\mathbf{c}(s{Q\over R})=\tilde{\mathbf{c}}(s{Q\over R})$ for $s=1,\cdots,R$.
The following theorem states that the barcode is parameterized by finite parameters $\mathbf{a}^{1},\cdots,\mathbf{a}^{s}$ in an explicit form.
Theorem 3.15.
For a piecewise linear approximation of a curve $\tilde{\mathbf{c}}$, we have
(16) $\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f})=\left\{J_{1}^{n_{1},\tilde{\mathbf{c}}}\bigcap\cdots\bigcap J_{N}^{n_{N},\tilde{\mathbf{c}}}:\sum\limits_{L=1}^{N}n_{L}=n\right\},$
where
$J_{L}^{n_{L},\tilde{\mathbf{c}}}=\begin{cases}\left(0,\infty\right),&\mbox{if }n_{L}=0\\ \left({2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\over\sqrt{N}a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}-{Q\over R}{\sum\limits_{i=1}^{{s_{b,L}}}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}\over a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}+{s_{b,L}}{Q\over R},{2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\over\sqrt{N}a_{L}^{s_{d,L}}/\lVert\mathbf{a}^{s_{d,L}}\rVert}-{Q\over R}{\sum\limits_{i=1}^{{s_{d,L}}}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}\over a_{L}^{s_{d,L}}/\lVert\mathbf{a}^{s_{d,L}}\rVert}+{s_{d,L}}{Q\over R}\right],&\mbox{if }n_{L}=2k+1\\ \ \emptyset,&\mbox{otherwise}\end{cases},$
where $s_{b,L}$ satisfies $\tilde{c}_{L}\left((s_{b,L}-1){Q\over R}\right)\leq 2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\leq\tilde{c}_{L}\left(s_{b,L}{Q\over R}\right)$ and $s_{d,L}$ satisfies $\tilde{c}_{L}\left((s_{d,L}-1){Q\over R}\right)\leq 2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\leq\tilde{c}_{L}\left(s_{d,L}{Q\over R}\right)$.
Proof.
Since
$\tilde{c}_{L}^{-1}(x)={x-\sqrt{N}{Q\over R}\sum\limits_{i=1}^{s-1}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}+\sqrt{N}{(s-1)\over R}Q{a_{L}^{s}\over\lVert\mathbf{a}^{s}\rVert}\over\sqrt{N}a_{L}^{s}/\lVert\mathbf{a}^{s}\rVert},\quad\tilde{c}_{L}\left((s-1){Q\over R}\right)\leq x\leq\tilde{c}_{L}\left(s{Q\over R}\right),$
the birth time is
$\tilde{c}_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\right)$ $={2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)-\sqrt{N}{Q\over R}\sum\limits_{i=1}^{s_{b,L}-1}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}+\sqrt{N}{(s_{b,L}-1)\over R}Q{a_{L}^{s_{b,L}}\over\lVert\mathbf{a}^{s_{b,L}}\rVert}\over\sqrt{N}a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}$
(17) $={2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\over\sqrt{N}a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}-{Q\over R}{\sum\limits_{i=1}^{{s_{b,L}}}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}\over a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}+{s_{b,L}}{Q\over R}.$
The death time can be shown similarly. ∎
Corollary 3.15.1.
The approximated barcode $\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f})$ converges to $\mathsf{bcd}_{n}^{\mathcal{R},\mathbf{c}}(\Psi_{f})$ as $R\rightarrow\infty$.
Proof.
In (17),
${2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\over\sqrt{N}a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}-{Q\over R}{\sum\limits_{i=1}^{{s_{b,L}}}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}\over a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}+{s_{b,L}}{Q\over R}$
$=\tilde{c}_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\right)$ $\longrightarrow\ c_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\right)$
as $R\rightarrow\infty$, and similarly,
${2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\over\sqrt{N}a_{L}^{s_{d,L}}/\lVert\mathbf{a}^{s_{d,L}}\rVert}-{Q\over R}{\sum\limits_{i=1}^{{s_{d,L}}}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}\over a_{L}^{s_{d,L}}/\lVert\mathbf{a}^{s_{d,L}}\rVert}+{s_{d,L}}{Q\over R}$
$=\tilde{c}_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\right)$ $\longrightarrow\ c_{L}^{-1}\left(2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\right)$
as $R\rightarrow\infty$. Since $J_{L}^{n_{L},\tilde{\mathbf{c}}}$ in (16) converges to $J_{L}^{n_{L},\mathbf{c}}$ in (14), $\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f})$ converges to $\mathsf{bcd}_{n}^{\mathcal{R},\mathbf{c}}(\Psi_{f})$ as $R\rightarrow\infty$.
Notations are shared in Table 2 except for $B:\left(\mathbb{R}_{m,M}^{N}\right)^{R}\rightarrow\mathsf{Bar},(\mathbf{a}^{1},\cdots,\mathbf{a}^{R})\mapsto\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f})$. As in Lemma 3.2 and Remark 3.3, $B$ and $V$ are continuous. Therefore $\mathcal{L}(\mathbf{W},\mathbf{b},\mathbf{a}^{1},\cdots,\mathbf{a}^{R}):=V\circ\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f})\rightarrow V\circ\mathsf{bcd}_{n}^{\mathcal{R},\mathbf{c}}(\Psi_{f})=\mathcal{L}(\mathbf{W},\mathbf{b},\mathbf{c})$ as $R\rightarrow\infty$, and the filtration curve learning problem can be approximated as the following constrained optimization problem:
(18) $\operatorname*{arg\,min}_{\begin{subarray}{c}\mathbf{W},\mathbf{b},\\ \mathbf{a}^{1},\cdots,\mathbf{a}^{R}\in\mathbb{R}_{m,M}^{N}\end{subarray}}\mathcal{L}(\mathbf{W},\mathbf{b},\mathbf{a}^{1},\cdots,\mathbf{a}^{R}).$
Figure 9 schematically illustrates the filtration curve learning of EMPH. The only difference from the filtration ray learning is the number of learning parameters of directions $\mathbf{a}^{s}$. Therefore, similar to Lemma 3.6, it can be shown that $\tilde{B}$ is locally Lipschitz and definable. Additionally, the gradient of the loss function with respect to the direction vector $\mathbf{a}^{s}$ can be derived. Consequently, Corollary 3.15.2 can be obtained.
Figure 9. Schematic illustration of filtration curve learning in EMPH.
Corollary 3.15.2.
For a piecewise linear approximation of a curve $\mathbf{c}$, as defined in (15), by the following update rule, the sequences $\mathbf{W}_{k},\mathbf{b}_{k}$ and $\mathbf{a}^{s}_{k}$ in (6) converge to Clarke stationary points of $\mathcal{L}$,
$W^{[l]}_{k+1}$ $=W^{[l]}_{k}-\alpha_{k}{\partial\mathcal{L}\over\partial{W^{[l]}_{k}}},$
$b^{[l]}_{k+1}$ $=b^{[l]}_{k}-\alpha_{k}{\partial\mathcal{L}\over\partial{b^{[l]}_{k}}},$
$\mathbf{a}_{k+{1\over 2}}^{s}$ $=\mathbf{a}_{k}^{s}-\alpha_{k}{\partial\mathcal{L}\over\partial{\mathbf{a}_{k}^{s}}},$
$\mathbf{a}_{k+1}^{s}$ $=\mathsf{proj}_{\mathbb{R}^{N}_{m,M}}\left(\boldsymbol{\rho}\left(\mathbf{a}_{k+{1\over 2}}^{s}\right)\right),$
where
${\partial\mathcal{L}\over\partial W^{[l]}}$ $=\left(\delta^{[l]}\right)\left(g^{[l-1]}(z_{l-1})\right)^{T},$
${\partial\mathcal{L}\over\partial b^{[l]}}$ $=\left(\delta^{[l]}\right),$
${\partial\mathcal{L}\over\partial\mathbf{a}^{s}}$ $=\left(\delta^{[1]}W^{[1]}\right)\left({1\over\mathbf{I}_{\max}-\mathbf{I}_{\min}}\right)\left({\partial\mathbf{I}\over\partial\mathsf{b}}{\partial\mathsf{b}\over\partial\boldsymbol{\rho}^{s}}+{\partial\mathbf{I}\over\partial\mathsf{d}}{\partial\mathsf{d}\over\partial\boldsymbol{\rho}^{s}}\right)\left({1\over\lVert\mathbf{a}^{s}\rVert}\mathbf{1}-{1\over\lVert\mathbf{a}^{s}\rVert^{3}}\mathbf{a}^{s}\left(\mathbf{a}^{s}\right)^{T}\right).$
The notations are shared in Theorem 3.11 and $\boldsymbol{\rho}^{s}={\mathbf{a}^{s}\over\lVert\mathbf{a}^{s}\rVert}=(\rho_{1}^{s},\cdots,\rho_{N}^{s})$.
Each component of ${\partial\mathcal{L}\over\partial\mathbf{a}^{s}}$ is as follows: the derivatives ${\partial I_{v}\over\partial b_{j}}$ and ${\partial I_{v}\over\partial d_{j}}$ are the same as in (8) and (9), respectively.
Clarke subgradients
${\partial b_{j}\over\partial\rho_{L}^{s}}\in$ $\ {\partial\over\partial\rho_{L}^{s}}\left(\max\left\{b_{L,n_{L}}(\boldsymbol{\rho}^{1},\cdots,\boldsymbol{\rho}^{R}):\sum\limits_{L}n_{L}=n\right\}\right),$
${\partial d_{j}\over\partial\rho_{L}^{s}}\in$ $\ {\partial\over\partial\rho_{L}^{s}}\left(\min\left\{d_{L,n_{L}}(\boldsymbol{\rho}^{1},\cdots,\boldsymbol{\rho}^{R}):\sum\limits_{L}n_{L}=n\right\}\right),$
where $(b_{j},d_{j}]=J_{1}^{n_{1},\tilde{\mathbf{c}}}\bigcap\cdots\bigcap J_{N}^{n_{N},\tilde{\mathbf{c}}}\in\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f})$ in (16) and
$b_{L,n_{L}}(\mathbf{a}^{1},\cdots,\mathbf{a}^{R})=\begin{cases}{2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\over\sqrt{N}a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}-{Q\over R}{\sum\limits_{i=1}^{{s_{b,L}}}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}\over a_{L}^{s_{b,L}}/\lVert\mathbf{a}^{s_{b,L}}\rVert}+{s_{b,L}}{Q\over R},&\mbox{if }n_{L}=2k+1\\ 0,&\mbox{otherwise }\end{cases},$
$d_{L,n_{L}}(\mathbf{a}^{1},\cdots,\mathbf{a}^{R})=\begin{cases}{2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\over\sqrt{N}a_{L}^{s_{d,L}}/\lVert\mathbf{a}^{s_{d,L}}\rVert}-{Q\over R}{\sum\limits_{i=1}^{{s_{d,L}}}{a_{L}^{i}\over\lVert\mathbf{a}^{i}\rVert}\over a_{L}^{s_{d,L}}/\lVert\mathbf{a}^{s_{d,L}}\rVert}+{s_{d,L}}{Q\over R},&\mbox{if }n_{L}=2k+1\\ \infty,&\mbox{otherwise }\end{cases}.$
Here, $s_{b,L}$ satisfies $\tilde{c}_{L}\left((s_{b,L}-1){Q\over R}\right)\leq 2r_{L}^{f}\sin\left(\pi{k\over 2k+1}\right)\leq\tilde{c}_{L}\left(s_{b,L}{Q\over R}\right)$ and $s_{d,L}$ satisfies $\tilde{c}_{L}\left((s_{d,L}-1){Q\over R}\right)\leq 2r_{L}^{f}\sin\left(\pi{k+1\over 2k+3}\right)\leq\tilde{c}_{L}\left(s_{d,L}{Q\over R}\right)$.
In particular, if we consider a one-dimensional barcode $(n=1)$, $j$ can be indexed such that $(b_{j},d_{j}]=\biggl{(}0,\sqrt{{3\over N}}r_{j}^{f}{1\over\rho_{j}^{s_{d,j}}}-{Q\over R}{\sum\limits_{i=1}^{s_{d,j}}\rho_{j}^{i}\over\rho_{j}^{s_{d,j}}}+s_{d,j}{Q\over R}\biggr{]}$. Then, we get
(19) ${\partial b_{j}\over\partial\rho_{L}^{s}}=$ $\ 0\quad\text{ for }1\leq j,L\leq N\text{ and }1\leq s\leq R,$
(20) ${\partial d_{j}\over\partial\rho_{L}^{s}}=$$\begin{cases}-{Q\over R}{1\over\rho_{L}^{s_{d,L}}},&\mbox{if }j=L\text{ and }s<s_{d,L},\\ -\sqrt{{3\over N}}r_{L}^{f}{1\over\left(\rho_{L}^{s_{d,L}}\right)^{2}}+{Q\over R}{\sum\limits_{i=1}^{s_{d,L}-1}\rho_{L}^{i}\over\left(\rho_{L}^{s_{d,L}}\right)^{2}},&\mbox{if }j=L\text{ and }s=s_{d,L},\\ 0,&\mbox{otherwise}\end{cases}.$
Proof.
Here, we only prove (19) and (20). Others can be proved similarly to the proof of Theorem 3.11. Since $s_{b,L}=0,b_{L,n_{L}}(\boldsymbol{\rho}^{1},\cdots,\boldsymbol{\rho}^{R})=0$ for every $L$ and $\sum\limits_{L}n_{L}=1$. Set $b_{j}=\max\bigl{\{}b_{j,n_{j}}(\boldsymbol{\rho}^{1},\cdots,\boldsymbol{\rho}^{R}):n_{j}=1\text{ and }n_{1},\cdots,n_{j-1},n_{j+1},\cdots,n_{N}=0\bigr{\}}=0$ and $d_{j}=\min\left\{d_{j,n_{j}}(\boldsymbol{\rho}^{1},\cdots,\boldsymbol{\rho}^{R}):n_{j}=1\text{ and }n_{1},\cdots,n_{j-1},n_{j+1},\cdots,n_{N}=0\right\}=\sqrt{{3\over N}}r_{j}^{f}{1\over\rho_{j}^{s_{d,j}}}-{Q\over R}{\sum\limits_{i=1}^{s_{d,j}}\rho_{j}^{i}\over\rho_{j}^{s_{d,j}}}+s_{d,j}{Q\over R}$. Differentiating these with respect to $\rho_{L}^{s}$ yields (19) and (20). ∎
Remark 3.16 (Comparison with vectorization of multi-parameter persistent homology).
The multi-parameter persistence image and multi-parameter persistence landscape are vectorization techniques derived from the rank invariant of multi-parameter persistent homology, as discussed in [22] and [30], respectively. Our method shares a common feature with these in considering the rank invariant (or fibered barcode). However, the difference lies in their examination of multiple lines in the multi-parameter space. Our approach focuses on identifying a single ‘optimal’ filtration curve (see Figure 10). This approach allows us to analyze the data more concisely and easily approach $n$-parameter persistence for $n\geq 2$. Additionally, we provide an analytic example of the multi-parameter persistence image and landscape in Appendix C.
$\Longrightarrow$
Figure 10. The multi-parameter persistence image and landscape consider multiple lines in the multi-parameter space, while our model pursues finding a single optimal filtration curve.
4. Numerical examples
In this section, we verify our proposed filtration learning model. In Example 4.1, we compare the computational efficiency of using the gradient formula versus not using the gradient formula in the backward step. In Examples 4.2 and 4.3, we demonstrate the efficiency of the proposed method of the filtration ray learning and filtration curve learning in synthetic data. In example 4.4, we compare the performances between our model filtration curve learning and other methods. These examples show that considering the optimal filtration curve is significant in the classification of time-series data.
The following Algorithm 1 summarizes our pipeline for solving the optimization problem in (18). In data analysis, time-series data are often given as discrete data. We assume the time-series data are evenly sampled with a length of $\mathfrak{n}$, and the Fourier coefficients are calculated using the Fast Fourier Transform.
1:  Input: $f_{1},\cdots,f_{m}:\left\{{2\pi i\over\mathfrak{n}}:0\leq i\leq\mathfrak{n}-1\right\}\rightarrow\mathbb{R}$ (time-series data)
2:  Variables: $E\in\mathbb{Z}_{\geq 0}$ (iterations), $N\in\mathbb{Z}_{\geq 0}$ (degree of truncated Fourier series), $R\in\mathbb{Z}_{\geq 0}$ (number of line segments), $n\in\mathbb{Z}_{\geq 0}$ (dimension of barcode) and $h\in\mathbb{Z}_{\geq 0}$ (number of hidden layers);
3:  Initialize the neural network and direction vectors: $\mathbf{W}_{0}:=\left(W^{[1]}_{0},\cdots,W^{[h]}_{0}\right),\mathbf{b}_{0}:=\left(b^{[1]}_{0},\cdots,b^{[h]}_{0}\right)$ and $\mathbf{a}^{s}_{0}=(1,\cdots,1)$ for $s=1,\cdots,R$.
4:  for $k=0,\cdots,E$ do
5:     for $j=1,\cdots,m$ do
6:        Calculate the Fourier transforms $\hat{f}_{j}$ using the fast Fourier transform.
7:        Calculate $\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f_{j}})$ using Theorem 3.15.
8:        Calculate persistence image of $\mathsf{bcd}_{n}^{\mathcal{R},\tilde{\mathbf{c}}}(\Psi_{f_{j}})$.
9:     end for
10:     Input the persistence image into the neural network.
11:     Apply backpropagation as follows:
$W^{[l]}_{k+1}$ $=W^{[l]}_{k}-\alpha_{k}{\partial\mathcal{L}\over\partial{W^{[l]}_{k}}},$
$b^{[l]}_{k+1}$ $=b^{[l]}_{k}-\alpha_{k}{\partial\mathcal{L}\over\partial{b^{[l]}_{k}}},$
$\mathbf{a}_{k+{1\over 2}}^{s}$ $=\mathbf{a}_{k}^{s}-\alpha_{k}{\partial\mathcal{L}\over\partial{\mathbf{a}_{k}^{s}}},$
$\mathbf{a}_{k+1}^{s}$ $=\mathsf{proj}_{\mathbb{R}^{N}_{m,M}}\left(\boldsymbol{\rho}\left(\mathbf{a}_{k+{1\over 2}}^{s}\right)\right).$
12:  end for
13:  return   Class prediction through the trained neural network
Algorithm 1 Filtration learning within EMPH (FL-EMPH)
Example 4.1.
In this example, we compare the computational efficiency of updating the direction vector $\mathbf{a}$ using the gradient formula versus not using the gradient formula in the backward step. We use a dataset consisting of 400 time-series data, each with a length of 80. We employ a 1-hidden layer neural network and vary the widths and resolution of the persistence image. All computations were performed using an i7-13700KF CPUs, 32GB of DDR5 RAM, and an RTX 4070 Ti GPUs.
We describe the models as follows:
  1. (1)
    PyTorch with direct update of the direction vector: The model utilizes the GPUs for both the forward and backward processes. It updates the direction vector $\mathbf{a}$ directly with the derivative ${\partial\mathcal{L}\over\partial\mathbf{a}}$ as shown in (7), while the neural network’s parameters are updated through PyTorch’s automatic differentiation.
  2. (2)
    Standard PyTorch: This model performs both forward and backward operations on the GPUs, with every parameter being updated via PyTorch’s automatic differentiation system.
  3. (3)
    Numpy and Scikit-learn: In this setup, both forward and backward operations are executed on the CPUs in Numpy and Scikit-learn. All parameters ($\mathbf{W}$, $\mathbf{b}$, and $\mathbf{a}$) are updated using the formula in (7).
Table 3 shows the computation times for 100 epochs of filtration learning. From these results, we can interpret that when the neural network is small, Numpy and Scikit-learn are superior to other methods, and when the neural network is large, PyTorch with direct update of the direction vector is superior. In conclusion, in both cases, using the gradient formula is beneficial in reducing computation times. The following examples were conducted on simple data, hence we carried out the experiments using Python code implemented with Numpy and Scikit-learn.
(Width of hidden layer, Resolution) PyTorch with direct update Standard PyTorch Numpy and Scikit-learn
$(100,10^{2})$ 19.1 s ± 19.6 ms 1min 8s ± 146 ms 5.88 s ± 101 ms
$(100,50^{2})$ 20.8 s ± 26.4 ms 1min 8s ± 86.3 ms 1min 42s ± 575 ms
$(100,100^{2})$ 23.1 s ± 19.4 ms 1min 10s ± 60.5 ms 5min 1s ± 2.56 s
$(1000,10^{2})$ 19.7 s ± 15.4 ms 1min 8s ± 172 ms 13 s ± 91.9 ms
$(1000,50^{2})$ 20.8 s ± 43.7 ms 1min 8s ± 67.9 ms 3min 18s ± 5.5 s
$(100,100^{2})$ 25.7 s ± 22 ms 1min 10s ± 55.8 ms 8min 13s ± 7.96 s
$(10000,10^{2})$ 20.1 s ± 38.3 ms 1min 8s ± 116 ms 1min 57s ± 2.17 s
$(10000,50^{2})$ 28.1 s ± 19.1 ms 1min 9s ± 56.8 ms 12min 15s ± 11.9 s
$(10000,100^{2})$ 59.2 s ± 108 ms 1min 13s ± 86.1 ms 45min 54s ± 7.52 s
Table 3. Computation times for 100 epochs of filtration ray learning.
Example 4.2.
Consider two time-series datasets, $g_{1}$ and $g_{2}$, each with a length of $36$, evenly sampled from $\cos t$ and $\cos 5t$, respectively, in the interval $[0,2\pi]$. We construct EMPH within the two-parameter space, which consists of $(\epsilon_{1},\epsilon_{5})$: a subspace of the frequency domain that includes the Fourier $1$-mode and $5$-mode. If we employ sliding window embedding or a diagonal ray for EMPH, we cannot distinguish between the two. Therefore, in such cases, our optimization method becomes useful when the objective is to distinguish between two time-series datasets. We set up the experiment as follows: We generated $100$ time-series data for each class and used $80\%$ of them for training, with the remainder used for testing. We use a 1-hidden layer with a width of 50, train for 10,000 epochs, set the learning rate to 0.001, use a bandwidth of 0.05 for the persistence image, and set the resolution of the persistence image to $10^{2}$.
$\left\{\begin{array}[]{l}g^{\epsilon}_{1}=\cos t+1\cdot\epsilon(0,1)\\ g^{\epsilon}_{2}=\cos 5t+1\cdot\epsilon(0,1)\end{array}\right.,$
where $\epsilon(0,1)$ represents the Gaussian standard normal errors.
We compare two methods: one where only the neural network is updated, resulting in an accuracy of 52.5 %, and the other, filtration ray learning, where both the neural network and the direction vector are updated, achieving an accuracy of 100 %. Figure 11 illustrates the loss and the transition of the direction vector under varying epochs.
Figure 11. Schematic illustration in Example 4.2. Left: Loss function in logarithmic scale. Red (accuracy: 52.5 %) represents the case where only the neural network is updated, while blue (accuracy: 100 %) represents the updating of both the neural network and the direction vector. Right: Variation of the direction vector under varying epochs.
Example 4.3.
Let $g_{11}=\cos t,g_{12}=\cos 2t,g_{2}=2\cos t$ and $g_{3}=2\cos 2t$. We construct EMPH within the two-parameter space, which consists of $(\epsilon_{1},\epsilon_{2})$: a subspace of the frequency domain that includes the Fourier $1$-mode and $2$-mode. To treat $g_{11}$ and $g_{12}$ as identical while differentiating between $g_{2}$ and $g_{3}$, simply selecting a ray falls short. Opting for the diagonal ray enables us to treat $g_{11}$ and $g_{12}$ as the same; yet, it fails to differentiate $g_{2}$ from $g_{3}$. On the other hand, opting for an alternative ray hinders our ability to treat $g_{11}$ and $g_{12}$ as equivalent. In Remark [14], a curved filtration approach was proposed. In this case, we address the classification problem by introducing a piecewise linear curve. The experimental setup is as follows: We generated 50 time-series data for each type.
$\left\{\begin{array}[]{l}g^{\epsilon}_{11}=\cos t+1\cdot\epsilon(0,1)\\ g^{\epsilon}_{12}=\cos 2t+1\cdot\epsilon(0,1)\\ g^{\epsilon}_{2}=2\cos t+1\cdot\epsilon(0,1)\\ g^{\epsilon}_{3}=2\cos 2t+1\cdot\epsilon(0,1)\end{array}\right.,$
where $\epsilon(0,1)$ represents the Gaussian standard normal errors.
The classification problem was divided into three classes, where the first class consists of $g^{\epsilon}_{11}$ and $g^{\epsilon}_{12}$, the second class consists of $g^{\epsilon}_{2}$ and the third class consists of $g^{\epsilon}_{3}$. For the experiment, 80% of the data were used for training and the rest for testing. We use a 1-hidden layer with a width 50, train for 10,000 epochs and set the resolution of the persistence image to $10^{2}$. 5-fold cross-validation was performed with learning rate options of $\left\{0.01,0.005\right\}$, bandwidth choices of $\left\{1,0.5,0.1\right\}$ and varying the number of line segments in filtration curve learning to $\left\{1,2,3\right\}$.
We compare three methods: one without filtration learning, which uses a fixed specific initial ray and updates only the neural network, resulting in an accuracy of 89.5 %. Another method, filtration ray learning, updates both the neural network and the direction vector, achieving an accuracy of 88.5 %. The third method, filtration curve learning, updates both the neural network and the filtration curve, achieving an accuracy of 91.5 %. The results represent the mean accuracy across 5 experiments. Figure 12 illustrates the loss and the transition of the filtration ray and filtration curve under varying epochs.
Figure 12. Schematic illustration in Example 4.3. Top left: Loss function in logarithmic scale. Green represents the case where only the neural network is updated, with a fixed curve $c(t)=(t,t)$ for $0\leq t\leq\sqrt{3}$ and $c(t)=(\sqrt{3},\sqrt{3})+\left(\sqrt{{3\over 2}}(t-\sqrt{3}),\sqrt{{1\over 2}}(t-\sqrt{3})\right)$ for $t\geq\sqrt{3}$. Red represents the case where both the neural network and filtration ray are updated, while blue represents the updating of both the neural network and the three direction vectors $\mathbf{a}^{1},\mathbf{a}^{2}$ and $\mathbf{a}^{3}$ (piecewise linear curve). Top right: Variation of the components of the direction vector. Bottom: Variation of the piecewise linear filtration curve under varying epochs.
Example 4.4.
In this example, we implement our theoretical approach on benchmarking data and evaluate its performance against other methods. Six datasets were selected from the UCR Time Series Classification Archive (https://www.cs.ucr.edu/~eamonn/time_series_data_2018/, [31] ) for this purpose. Our proposed model is Filtration learning within EMPH (FL-EMPH). We use a 2-hidden layer with a width 50, 10, respectively, train for 10,000 epochs and set the resolution of the persistence image to $10^{2}$. 5-fold cross-validation was performed with learning rate options of $\left\{0.01,0.005\right\}$, bandwidth choices of $\left\{1,0.5,0.1\right\}$ and varying the number of line segments in filtration curve learning to $\left\{1,2,3,5\right\}$. Table 4 shows the classification accuracy. The results represent the mean accuracy across 5 experiments. We compare our method (FL-EMPH) not only with TDA methods but also with other approaches. These include
  1. (1)
    EMPH: The EMPH model employs topological featurization as in EMPH and updates only the neural network, excluding filtration learning
  2. (2)
    MPI: Multi-parameter Persistence Image, [22]
  3. (3)
    PI: Persistence Image, [2]
  4. (4)
    MPL: Multi-parameter Persistence Landscape, [30]
  5. (5)
    PL: Persistence Landscape, [1]
  6. (6)
    MPK: Multi-parameter Persistence Kernel, [32]
  7. (7)
    PSSK: Persistence Scale Space Kernel, [33].
  8. (8)
    B1: Euclidean nearest neighbor with 1-nearest neighbor classification.
  9. (9)
    B2: Dynamic time warping 1-nearest neighbor classification.
  10. (10)
    B3: Constant window width 1-nearest neighbor classification.
B1, B2 and B3 are introduced in [31].
Table 4. Classification accuracy (%). The results from MPI to B3 are cited in [22].
From the result in Table 4, it is evident that filtration learning effectively enhances the accuracy of classification. Moreover, when comparing performance with other methods, the FL-EMPH approach further motivates its application to more complex datasets.
5. Conclusion
In this paper, we proposed an optimization problem to find an optimal filtration curve within EMPH for the classification of time-series data. Filtration learning involves updating filtration, calculating the persistence image and updating neural network for each epoch; this process is computationally extremely heavy. In our case, from the exact formula of the barcode, we can deduce the gradient ${\partial\mathcal{L}\over\partial\mathbf{a}}$. Therefore, we can directly update $\mathbf{a}$ without automatic differentiation. Our approach effectively reduces the computational complexity of forward and backward steps, thereby facilitating efficient filtration learning. Moreover, our experiments demonstrate that focusing on finding an optimal filtration curve is meaningful in terms of classification accuracy, and generalizing in a higher-dimensional parameter space is easier than considering multiple lines within the parameter space. Meanwhile, as the computational costs remain high, incorporating our model to reduce the dimensionality of the multi-parameter space seems to be a considerable approach. Our model can enhance its performance by incorporating numerous techniques from deep learning (for example, minibatch learning, dropout, early stopping), and we will apply this to real-world data. Additionally, while we have so far conducted experiments only with one-dimensional persistent homology, we have plans to extend our experiment to include higher-dimensional persistent homology.
Acknowledgement This work was supported by NRF of Korea under the grant number 2021R1A2C3009648 and POSTECH Basic Science Research Institute under the NRF grant number 2021R1A6A1A1004294412. This work is also supported partially by NRF grant by the Korea government (MSIT) (RS-2023-00219980).
Appendix A o-minimal structure
This section is based on [34, 35, 36, 26].
The motivation of o-minimal structure is consider the class of nice subsets of $\mathbb{R}^{n}$ and class of nice functions between the nice set. Nice means the class closed under elementary logical (unions, intersections, complement), geometric and topological operation (closure, cartesian products, project into lower dimensional subspace). One of the example of the class of nice set is semialgebraic sets. In the name o-minimal, ‘o’ stands for order.
Definition A.1 (Semialgebraic set, Semialgebraic function).
A subset $X\subseteq\mathbb{R}^{n}$ is semialgebraic if $X$ is a finite union of $\left\{x\in\mathbb{R}^{n}:P(x)=0\text{ and }Q_{1}(x),\cdots,Q_{l}(x)>0\right\}$, where $l\in\mathbb{Z}_{\geq 0}$ and $P,Q_{1},\cdots,Q_{l}\in\mathbb{R}[x_{1},\cdots,x_{n}]$. For semialgebraic sets $A\subseteq\mathbb{R}^{m}$ and $B\subseteq\mathbb{R}^{n}$, a function $f:A\rightarrow B$ is called semialgebraic if its graph is a semialgebraic subset of $\mathbb{R}^{m}\times\mathbb{R}^{n}$.
Boolean algebra is an axiomatic extension of power set.
Definition A.2 (Boolean algebra, [37]).
Boolean algebra is a nonempty set $\mathcal{A}$ with binary operations $\lor$ and $\land$ on $\mathcal{A}$ and a unary operation and two distinguished elements 00 and $1$ in $\mathcal{A}$ satisfying suitable operation rules such as in a power set of nonempty set.
$\mathcal{A}$longleftrightarrow $\mathcal{P}(X)$
[…]
$1$longleftrightarrow $X$
Table 5. Correspondence between Boolean algebra and a power set of nonempty set.
o-minimal structure is an axiomatic extension of semialgebraic set.
Definition A.3.
o-minimal structure on the real number field $(\mathbb{R},+,\cdot)$ is a sequence of $\mathcal{D}=(\mathcal{D}_{n})_{n\in\mathbb{N}}$ such that
  1. (1)
    $\mathcal{D}_{n}$ is a boolean algebra of subsets of $\mathbb{R}^{n}$. i.e.,$\emptyset,\mathbb{R}^{n}\in\mathcal{D}_{n}$, and if $A,B\in\mathcal{D}_{n}$, then $A\cup B\in\mathcal{D}_{n}$ and $\mathbb{R}^{n}\setminus A\in\mathcal{D}_{n}$.
  2. (2)
    If $A\in\mathcal{D}_{n}$, then $A\times\mathbb{R}$ and ×$\mathbb{R}\times A\in\mathcal{D}_{n+1}$.
  3. (3)
    If $A\in\mathcal{D}_{n+1}$, then $\pi(A)\in\mathcal{D}_{n}$, where $\pi:\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}^{n}$ is the projection on the first $n$ coordinates.
  4. (4)
    $\mathcal{D}_{n}$ contains $\left\{x\in\mathbb{R}^{n}:p(x)=0\right\}$ for every polynomial $p\in\mathbb{R}[x_{1},\cdots,x_{n}]$.
  5. (5)
    The elements of $\mathcal{D}_{1}$ are exactly the finite unions of intervals and points. (This is called o-minimal constraint)
Definition A.4.
We say $A\subset\mathbb{R}^{m}$ is definable set if $A\in\mathcal{D}_{n}$. For a definable set $A$, a function $f:A\rightarrow\mathbb{R}^{n}$ is called definable function if its graph is definable set in $\mathbb{R}^{m+n}$.
Proposition A.5.
Suppose $A\subseteq\mathbb{R}^{m}$ is definable set and $f:A\rightarrow\mathbb{R}^{n}$ is a definable function. Then the followings are hold.
  1. (1)
    Let $A\in\mathcal{D}_{m}$ and $B\in\mathcal{D}_{n}$. Then closure, interior of $A$ is also definable, i.e., $cl(A),int(A)\in\mathcal{D}_{m}$. Additionally, $A\times B\in\mathcal{D}_{m+n}$ also holds [34].
  2. (2)
    Sum, product, composition, restriction and inverse of definable functions is also definable. Image of definable function is a definable set. Moreover, $f=(f_{1},\cdots,f_{n})$ is definable if and only if $f_{i}$ are definable [34].
  3. (3)
    Each definable set $A\in\mathbb{R}^{m}$ has a finite partition $A=C_{1}\cup\cdots\cup C_{k}$ into $C^{1}$-cells and $f|_{\mathcal{C}_{i}}$ are $C^{1}$ [34].
  4. (4)
    There is an o-minimal structure that contains the graph of the exponential function and semialgebraic sets [26, 38].
Proposition A.6 ([39]).
Let $f:X\rightarrow\mathbb{R}$ be a semialgebraic function. If $f(x)\neq 0$ for every $x\in X$, then ${1\over f}$ is semialgebraic, and if $f\geq 0$, then $\sqrt{f}$ is also a semialgebraic function.
Proposition A.7 ([39]).
If $f$ and $g$ are semi-algebraic, then $\min\left\{f,g\right\}$ is also semi-algebraic.
Proposition A.8 ([36]).
Every semialgebraic set is definable in any o-minimal structure.
Appendix B Propositions for showing $\mathcal{L}$ is locally Lipschitz
Proposition B.1.
For locally Lipschitz functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ and $g:\mathbb{R}^{m}\rightarrow\mathbb{R}$, $g\circ f$ is also locally Lipschitz.
Proof.
Let $a=f(x)$. Since $g$ is locally Lipschitz, there is a neighborhood $U_{a}$ of $a$ such that $g|_{U_{a}}$ is Lipschitz. i.e., there is $L_{1}\in\mathbb{R}$ such that for every $b\in U_{a}$, $\left|g(a)-g(b)\right|\leq L_{1}\left|a-b\right|$. Given that $f$ is continuous, there is a neighborhood $V_{x}$ of $x$ such that $f(V_{x})\subseteq U_{a}$. Since $f$ is locally Lipschitz, there is a $L_{2}\in\mathbb{R}$ and a neighborhood $W_{x}$ such that $\left|f(x)-f(y)\right|\leq L_{2}\left|x-y\right|$ for every $y\in W_{x}$. Consequently, for any $y\in V_{x}\bigcap W_{x}$, it follows that $\left|g(f(x))-g(f(y))\right|\leq L_{1}\left|f(x)-f(y)\right|\leq L_{1}L_{2}\left|x-y\right|$. ∎
Proposition B.2 (Exercise 4.2.10, [40]).
If $X$ is a locally compact metric space, then $f$ is locally Lipschitz if and only if it is Lipschitz continuous on every compact subset of $X$.
Proposition B.3.
If $f_{1},f_{2}$ are locally Lipschitz, then $\min\left\{f_{1},f_{2}\right\}$ and $\max\left\{f_{1},f_{2}\right\}$ are locally Lipschitz.
Proof.
Since the sum of locally Lipschitz is locally Lipschitz and $\left|f_{1}-f_{2}\right|$ is locally Lipschitz, $\min\left\{f_{1},f_{2}\right\}={f_{1}+f_{2}-\left|f_{1}-f_{2}\right|\over 2}$ is locally Lipschitz. Maximum function case can be proved similarly. ∎
Appendix C Vectorization of Multi-parameter persistent homology
In this section, we review the vectorization of multi-parameter persistent homology, using multi-parameter persistence image and landscape as examples, to compare our method. The following definitions are based on the Python code  [^1] https://github.com/MathieuCarriere/multipers, [22]  [^2] https://github.com/OliverVipond/Multiparameter_Persistence_Landscapes, [30].
Definition C.1 (Two-parameter persistence image, [22]).
Consider a grid $\mathcal{G}=\left\{\mathbf{g}_{i,j}\right\}_{0\leq i,j\leq n-1}$, where each $\mathbf{g}_{i,j}$ is defined as $\left(a+i\frac{b-a}{n-1},c+j\frac{d-c}{n-1}\right)$. Denote $\ell(x,v)$ as a ray passing through $x$ with direction vector $v$. Let $L=\biggl{\{}\ell\left((a+i\delta,c),\left({\sqrt{2}\over 2},{\sqrt{2}\over 2}\right)\right):0\leq i\leq\lfloor{\#\text{rays}\over 2}\rfloor\biggr{\}}\bigcup\left\{\ell\left((a,c+j\delta),\left({\sqrt{2}\over 2},{\sqrt{2}\over 2}\right)\right):0\leq j\leq\lfloor{\#\text{rays}\over 2}\rfloor\right\}$ be a collection of rays, where $\delta={(b-a)+(d-c)\over\#\text{rays}}$. For a Two-parameter persistent homology $\mathsf{MP}$, we construct an $n^{2}$-dimensional vector $\mathcal{I}=(\mathcal{I}_{i,j})_{0\leq i,j\leq n-1}$ where
$\mathcal{I}_{i,j}=\sum\limits_{I\in D_{L}(\mathsf{MP})}w_{1}(I)\left(w_{2}(\ell^{*})\exp\left(-{\underset{\ell\in I}{\min}\lVert\mathbf{g}_{i,j},\ell\rVert^{2}\over\sigma^{2}}\right)\right),$
and in this context, $D_{L}(\mathsf{MP})$ is the vineyard decomposition of the fibered barcodes $\left\{\mathsf{MP}^{\ell_{i}}\right\}$, $\lVert\mathbf{g}_{i,j},\ell\rVert$ is defined as $\underset{x\in\ell}{\min}\lVert\mathbf{g}_{i,j}-x\rVert$. Here, $\ell^{*}$ denotes the ray for which $\underset{\ell\in I}{\min}\lVert\mathbf{g}_{i,j},\ell\rVert^{2}$ is achieved, $A(I)$ represents the area of the convex hull formed by the endpoints of all bars in $I$, the function $w_{1}:D_{L}(\mathsf{MP})\rightarrow\mathbb{R}$ is given by $w_{1}(I)=\left(\frac{A(I)}{(b-a)(d-c)}\right)^{q}$, and $w_{2}(\ell)$ is $\min(e^{\ell}_{1},e^{\ell}_{2})$, where $(e^{\ell}_{1},e^{\ell}_{2})$ is the direction vector of $\ell$.
Definition C.2 (Two-parameter persistence landscape, [30]).
All notations are shared in Definition C.1. Denote rays by
$\ell_{i}=\begin{cases}\ell\left((a-i{b-a\over n-1},c),\left(1,1\right)\right),&\mbox{if }-(n-1)\leq i\leq 0\\ \ell\Bigl{(}(a,c+i{d-c\over n-1}),\left(1,1\right)\Bigr{)},&\mbox{if }0<i\leq n-1\end{cases},$
and endpoint of rays $l_{i}$ by
$\mathbf{o}_{i}=\begin{cases}(a-i{b-a\over n-1},c),&\mbox{if }-(n-1)\leq i\leq 0\\ (a,c+i{d-c\over n-1}),&\mbox{if }0<i\leq n-1\end{cases}.$
Define a $k$-th two-parameter persistence landscape $\mathcal{S}=(\mathcal{S}_{i,j})_{0\leq i,j\leq n-1}\in\mathbb{R}^{n^{2}}$, where
$\mathcal{S}_{i,j}=k\text{-th largest value of }\bigcup\limits_{s}\left\{\max\left\{0,\min\left\{{\lVert\mathbf{g}_{i,j}-\mathbf{o}_{j-i}\rVert-b_{s}\over\sqrt{2}},{d_{s}-\lVert\mathbf{g}_{i,j}-\mathbf{o}_{j-i}\rVert\over\sqrt{2}}\right\}\right\}\right\},$
where $(b_{s},d_{s}]\in\mathsf{bcd}^{\ell_{j-i}}_{*}(\mathsf{MP})$.
Although Definition C.1 and C.2 may seem complicated, we have prepared the following simple example.
Example C.3.
Let $\mathcal{G}$ be a grid with $a=c=0$,$b=d=2$ and $n=2$ in Definition C.1. Take the graph $X=\left\{a,b,c,d,ac,bc\right\}$ and define a two-parameter filtration $F:X\rightarrow\mathbb{R}^{2}$ with $F(a)=(0,1),F(b)=(1,0),F(c)=(1,1),F(d)=(1,1),F(ac)=(1,1)$, and $F(bc)=(1,1)$. Define a two-parameter persistent homology $\mathsf{MP}$ as $\mathsf{MP}(r_{1},r_{2}):=H_{0}\left(F^{-1}(r_{1},r_{2})\right)$.
  1. (1)
    (Two-parameter persistence image) Set $q=1$ and $\sigma=1$ in Definition C.1. The Vineyard Decomposition $D_{L}(\mathsf{MP})=\left\{I_{1},I_{2}\right\}$, where $I_{1}$ consists of three blue lines and $I_{2}$ consists of a single red line as shown in Figure 13. For the components of the two-parameter persistence image, the calculations are as follows: $A(I_{1})={20\over 9},A(I_{2})=0,w_{1}(I_{1})={5\over 9},w_{1}(I_{2})=0$, and $w_{2}(\ell^{*})={\sqrt{2}\over 2}$.
    $\mathcal{I}_{00}$ $={5\over 9}\cdot{\sqrt{2}\over 2}\cdot\exp\left(-{16\over 9}\right)\approx 0.07,$
    $\mathcal{I}_{01}$ $={5\over 9}\cdot{\sqrt{2}\over 2}\cdot\exp\left(-{2\over 9}\right)\approx 0.31,$
    $\mathcal{I}_{10}$ $={5\over 9}\cdot{\sqrt{2}\over 2}\cdot\exp\left(-{2\over 9}\right)\approx 0.31,$
    $\mathcal{I}_{11}$ $={5\over 9}\cdot{\sqrt{2}\over 2}\cdot\exp\left(0\right)\approx 0.39.$
    Figure 13. Supplementary of two-parameter persistence image.
  2. (2)
    (Two-parameter persistence landscape) Set $k=1$ in Definition C.2. Then $\ell_{-1}=\ell\left((2,0),(1,1)\right),\ell_{0}=\ell\left((0,0),(1,1)\right)$ and $\ell_{1}=\ell\left((0,2),(1,1)\right)$. The fibered barcodes are $\mathsf{bcd}^{\ell_{-1}}_{0}(\mathsf{MP})=\left\{(0,\infty),(\sqrt{2},\infty)\right\},$ $\mathsf{bcd}^{\ell_{0}}_{0}(\mathsf{MP})=\left\{(\sqrt{2},\infty)_{(2)}\right\}$ and $\mathsf{bcd}^{\ell_{1}}_{0}(\mathsf{MP})=\left\{(0,\infty),(\sqrt{2},\infty)\right\}$. The distances $\lVert\mathbf{g}_{0,0}-\mathbf{o}_{0}\rVert=\lVert\mathbf{g}_{1,0}-\mathbf{o}_{-1}\rVert=\lVert\mathbf{g}_{0,1}-\mathbf{o}_{1}\rVert=0$ and $\lVert\mathbf{g}_{1,1}-\mathbf{o}_{0}\rVert=2\sqrt{2}$.
[…]
    Figure 14. Supplementary of two-parameter persistence landscape.